### v1

# Unlock the Power of Frozen LLMs in Knowledge Graph Completion

Abstract
Classical knowledge graph completion (KGC) methods rely solely on structural information, struggling with the inherent sparsity of knowledge graphs (KGs). Large Language Models (LLMs) learn extensive knowledge from large corpora with powerful context modeling, which is ideal for mitigating the limitations of previous methods. Directly fine-tuning LLMs offers great capability but comes at the cost of huge time and memory consumption, while utilizing frozen LLMs yields suboptimal results. In this work, we aim to leverage LLMs for KGC effectively and efficiently. We capture the context-aware hidden states of knowledge triples by employing prompts to stimulate the intermediate layers of LLMs. We then train a data-efficient classifier on these hidden states to harness the inherent capabilities of frozen LLMs in KGC. We also generate entity descriptions with subgraph sampling on KGs, reducing the ambiguity of triplets and enriching the knowledge representation. Extensive experiments on standard benchmarks showcase the efficiency and effectiveness of our approach. We outperform classical KGC methods on most datasets and match the performance of fine-tuned LLMs. Additionally, compared to fine-tuned LLMs, we boost GPU memory efficiency by 188×188$\times$188 × and speed up training+inference by 13.48×13.48$\times$13.48 ×.
1 Introduction
Knowledge graph completion has become a crucial endeavor for the effective application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b). Classical knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which suffer from the sparseness of KGs (Yao et al., 2019). Leveraging the powerful contextual modeling capabilities and extensive stored world knowledge of LLMs can mitigate this challenge by utilizing the rich semantics within knowledge graphs, thereby enhancing the performance of knowledge graph completion (Wei et al., 2023).
These approaches generally fall into two types: using frozen LLMs and employing fine-tuned LLMs. The first type leverages frozen language models (e.g., LLaMA and GPT-4) with prompt techniques, such as in-context learning, which facilitate the completion of knowledge graphs directly in natural language outputs (Wei et al., 2023; Yao et al., 2023), offering an efficient solution that utilizes the pre-trained capabilities of these models without the need for further training. The second type involves fine-tuning the language models using a range of supervised techniques, including instruction tuning, to tailor the models specifically for KGC tasks. This method involves fine-tuning language models, such as LLaMA and T5, through the use of training samples (Saxena et al., 2022; Yao et al., 2023). Although this can lead to better performance (Zhang et al., 2023; Yao et al., 2023), it also requires significant increases in time and resource expenditure. A natural question is: How can we utilize LLMs to accomplish KGC tasks with both efficacy and efficiency?
To answer this question, we build upon the efficiency of frozen LLMs and explore the reasons behind the underperformance of frozen LLMs compared to their fine-tuned counterparts. First, while LLMs contain vast amounts of world knowledge encoded within their parameters, accessing the knowledge and decoding it into responses effectively through frozen models can be challenging (Wang et al., 2023). It is well known that LLMs suffer from hallucinations (Zhao et al., 2023), which makes them unreliable for achieving high performance in KGC tasks. Fine-tuning directly modifies the models’ weights through the training samples to improve the retrieval and application of relevant world knowledge, thus leading to better performance (Zhou et al., 2024). Second, KGs often involve complex relationships and ambiguous entity representations (Li et al., 2024), which can be difficult for a frozen LLM to handle accurately without specific tuning. Fine-tuned models can be explicitly trained to interpret and disambiguate such complexities on the basis of the training data provided, linking entities in the KG with the LLM.
To overcome the limitations of frozen LLMs in KGC tasks, we have introduced probing techniques utilizing stimulation prompts, complemented by detailed descriptions of entities, as shown in Figure 1. These methods are designed to effectively extract and leverage the latent knowledge embedded within the models’ parameters, and minimize ambiguous entity representations, thereby enhancing its application in KGC without the need for extensive retraining. Specifically, our research concentrates on the core issues of knowledge graph completion, which are triple classification and relation prediction. We stimulate the intermediate layers of LLMs to capture the context-aware hidden states of knowledge triples. Moreover, we improve the precision of knowledge representation by aligning KG entities with the LLMs, reducing the ambiguity of knowledge triples through the generation of entity descriptions via subgraph sampling from KGs. We train a data-efficient classifier on these obtained hidden states suppervised by the labels, to fully harness the inherent KGC capabilities of the frozen LLMs.
Figure 1: The overall architecture of our proposed probing method for the triple classification. Constructing positive and negative sample pairs, followed by stimulating the middle layers of the language model to obtain hidden states for completing the KGC task. Entity descriptions are generated by the subgraph entity description generator. The only component that needs training in the entire architecture is the data-efficient classifier used to classify hidden states, which can be MLP, logistic regression, etc.
Our contributions can be summarized as follows.
  • We introduce a novel probing method tailored for KGC tasks, designed to stimulate language models to verify knowledge graph triples directly. We further generate knowledge graph entity descriptions through the use of subgraphs, which are then seamlessly integrated into our probing techniques.
  • Our approach employs LLMs with both effectiveness and efficiency to achieve KGC. By utilizing only a subset of the dataset, our method achieves a relative performance improvement of up to 2.7% compared to fine-tuning the LLM on the entire training set. Additionally, it significantly reduces GPU memory usage by a factor of 188 during the training phase and descreases time consumption by a factor of 13.48, all while maintaining comparable performance. [^1] For a detailed comparison of efficiency, please refer to Table 8 in the Appendix B.3.
  • We conduct extensive experiments on six standard knowledge graph datasets to benchmark our method against multiple baseline approaches. We investigate the versatility of our method with various language models, assessing its efficacy and data efficiency through rigorous experimental validation.
2 Method
Our method’s design stems from an analysis of the two primary reasons for the underperformance of frozen LLMs compared to fine-tuned counterparts. First, LLM may not always trigger the depth of knowledge needed and reveal the right responses. We demonstrate in Section 2.2 how to achieve this for the frozen LLM using probing techniques. Second, KGs often involve ambiguous entity representations, which can be difficult for LLM to handle accurately. In Section 2.3 we show how to achieve this with the generation of entity descriptions.
2.1 Formulation and Notations
A knowledge graph can be represented as a set of triples $G=\{(h,r,t)\}$, where $E$ and $R$ denote the set of entities and relations in G, respectively, $h\in E$ is a head entity, $t\in E$ is a tail entity, we collectively refer to the head entity and tail entity as $e$, and $r\in R$ represents the relation between them. $D$ is the set of descriptions for each entity and relation. We denote $D(e)$, $D(r)$ as the textual description of each entity and each relation.
We focus on triple classification and relation prediction. Triple classification involves evaluating the plausibility of a given knowledge triple, to determine whether the triple is true or not and relation prediction predicts the missing relation in a triple.
2.2 Stimulating the Intermediate Layers
Knowledge graphs consist of numerous triples, and the format of these triples is challenging for the LLM to understand, which may mislead the LLM to retrieve irrelevant knowledge and reveal the wrong responses. To minimize the influence of task-unrelated information and enhance the performance of LLMs in KGC, we leverage the internal judgment mechanisms of the language model. Our method involves strategically triggering the necessary depth of knowledge embedded within the LLM to elicit the correct responses required for accurate KGC. By focusing on extracting classification results directly from the model’s internal processing, we ensure that the responses are both relevant and precise, thereby improving the effectiveness of the KGC task.
Since different layers of the language model have different roles (Azaria and Mitchell, 2023), we determine the most suitable layer for the task by experimenting with the validation set. First, we construct prompt templates to stimulate the model to generate the required classification information in the intermediate layers. In the triple classification, our goal is to determine whether the triples in the knowledge graph are true. Therefore, we use the following prompt template, $PT_{1}$, to create positive and negative sample pairs, as shown in Figure 2. Notice that the prompt does not describe positive or negative information for triples.
Then, we use existing triples in the knowledge graph to construct positive samples $S^{+}$, and employ the negative sampling technique (Lv et al., 2022) to obtain negative samples $S^{-}$. Finally, we obtain $N$ training data samples. For each training text, we input it into the language model to obtain the representation of each text in the intermediate layers. We can obtain information about the plausibility of triples from the hidden state corresponding to the token immediately preceding the model prediction, which is the last token of the input text (Zou et al., 2023). Specifically, we capture plausibility information for positive samples $s^{+}\in S^{+}$ by:
$v\left(s^{+}\right)=\{M_{1-i}\left(s^{+}\right)[-1],0<i<L\},$ (1)
where $M$ represents the language model, $M_{1-i}$ represents the first $i$ layers of the model, $[-1]$ refers to the hidden state corresponding to the last token, $L$ denotes the number of layers in the model. The dimension of $M_{1-i}(s^{+})[-1]$ is the embedding dimension of the language model. Negative samples follow the same principle.
At last, we train a classification model using $v(s^{+})$ and $v(s^{-})$ to capture plausibility information. The goal is to learn a boundary that can separate the hidden states of the intermediate layers of the language model corresponding to positive samples (which are true) from those corresponding to negative samples (which are false). We can use classification models such as multilayer perceptron (MLP), SVM, etc., for classification training. Taking logistic regression to determine the plausibility of samples as an example:
$P\left(Y=\mathbf{1}|x\right)=\frac{1}{1+e^{-\left(w^{T}x+b\right)}},$ (2)
where x represents $v(s)[l]$, $l$ is the layer index, $w$ and $b$ are weights and bias to be learned. For relation prediction, we simply transform the binary classification into multi-class classification. By utilizing the prompt in Figure 7, we stimulate the LLM to generate hidden states that capture the characteristics of different relation types in the KG. Details can be found in the Appendix B.1.
The task-unrelated information does not affect the classification results. The reason is that the positive samples $S^{+}$ and negative samples $S^{-}$ are constructed using the same prompt template. The classifier only needs to capture the plausibility of the triples in $v(S^{+})$ and $v(S^{-})$, that is, whether the triple is true or not. The advantage of using this method for KGC is that it can ignore the impact of the task-unrelated information and reveal the right responses for accurate KGC.
Figure 2: Prompt templates to stimulate the LLM for triple classification.
2.3 Subgraph Entity Description Generator
The semantics of entities in the knowledge graph are highly complex; it is difficult for the language model to understand the entity based solely on the entity’s name. Therefore, we inform the language model about the meaning of each entity in the input. The meaning of each entity in KG is generated by its one-hop subgraph.
Specifically, we extend the prompt template constructed in Section 2.2, aiming to assist the language model in aligning internal entities with entities in the knowledge graph. For triple classification, since we employ the probing method to extract KGC features from the samples, we need to ensure that the language model can be stimulated to produce effective hidden states that can be used by the classifier to predict the plausibility of triples. Therefore, we construct a new prompt $PT_{2}$, as shown in Figure 2. By introducing the $D(e)$, LLMs are able to generate more precise and distinguishable hidden states, thereby improving the overall classification performance.
To address incomplete entity descriptions within certain datasets, we have devised a method to generate $D(e)$. We construct $D(e)$ in two ways. The first approach involves a straightforward method where we concatenate all the triples within the one-hop subgraph surrounding the entity $e$, thereby forming $D(e)$. However, this direct concatenation can sometimes lead to ineffective or misleading stimuli for the LLM, as the structure and content of these subgraph triples may significantly differ from the text seen during the LLM’s pre-training phase. This divergence can adversely affect the model’s response accuracy. To mitigate this issue, our second approach leverages a language model to transform these triples into a format that is more comprehensible to the LLM. By rephrasing the triples from the one-hop subgraph around the entity $e$ into natural language text, we produce a more refined and model-friendly version of $D(e)$:
$D(e)=\operatorname{concat}\{\operatorname{transform}(\{h_{1}\}\{r_{1}\}\{t_{1}\}),$ (3)
$(h_{1},r_{1},t_{1})\in\operatorname{subgraph}(e)\},$
where $\operatorname{subgraph(\cdot)}$ is a function that retrieves the set of all triples in the one-hop subgraph where the entity resides; $\operatorname{transform(\cdot)}$ is a function that converts the triple into sentence; $\operatorname{concat(\cdot)}$ is a function to process triple statements. For the first approach, $\operatorname{concat(\cdot)}$ is simply a string concatenation function that joins each transformed triple statement with a separator. For the second approach, we need to utilize LLM with the in-context learning technique (Brown et al., 2020) to generate descriptions of entities based on the triples in their subgraph. Please refer to Appendix A.1 for details.
By generating long entity descriptions using the subgraph description generator, we align the internal understanding of entities in the LLM with the entity meanings in the KG, which is similar to the entity linking process in knowledge graphs. In this way, the LLM will better understand the semantics of the entities, and improve the KGC performance.
3 Experiments
In the experiments, our objective is to answer the following research questions (RQs):
  • RQ1: Can frozen LLMs with the proposed probing methods effectively achieve performance on par with their fine-tuned counterparts?
  • RQ2: How does our method perform across different language models and classification models?
  • RQ3: How about data efficiency and compute efficiency of our proposed method?
3.1 Datasets
Dataset |E| |R| # Train # Valid # Test
FB13 75,043 13 316,232 5,908 23,733
WN11 38,696 11 112,581 2,609 10,544
FB15K-237N 13,104 93 87,282 14,082 16,452
WN18RR 40,943 11 86,835 3,034 3,134
UMLS 135 46 5,216 1,304 1,322
YAGO3-10 123,182 37 1,079,040 5,000 5,000
Table 1: Statistical information of datasets.
In our experiments, we employ five widely-used KG datasets for triple classification: FB13 (Socher et al., 2013), WN11 (Socher et al., 2013), FB15K-237N (Lv et al., 2022), WN18RR (Dettmers et al., 2018), and UMLS (Yao et al., 2019), as well as a commonly used KG dataset for relation prediction: YAGO3-10 (Dettmers et al., 2018).
For FB13, we use Wikipedia entity descriptions as input sentences. For FB15K-237N, the entity descriptions provided by (Xie et al., 2016) are used. Synset definitions from (Yao et al., 2019) served as entity descriptions for WN18RR. In the case of UMLS, entity descriptions provided by the dataset are used. In the following, we refer to the entity descriptions as non-generated entity descriptions.
3.2 Experimental Settings
To generate entity descriptions, for the first approach in Section 2.3, we concatenate all subgraph triples to generate descriptions. For the second approach, we employ GPT-3.5-turbo as the LLM. Due to cost considerations, we randomly select 4000 test examples from each dataset to generate entity descriptions. We mark the results of these methods with an asterisk (*). In the following text, we refer to these entity descriptions as generated entity descriptions(Tri)/(GPT).
We compare our method with four structure information-based methods: TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), ComplEx (Trouillon et al., 2016) and RotatE (Sun et al., 2019), as well as seven additional information-based methods: KG-LLAMA (Yao et al., 2023), KG-BERT (Yao et al., 2019), KG-T5 (Saxena et al., 2022), LLaMA-7b, Alpaca-7B, LLaMA-7B-ICL and Alpaca-7B-ICL (Zhang et al., 2023). KG-LLAMA, KG-BERT and KGT5 fine-tune the language models using all the training samples. LLaMA-7b and Alpaca-7B using the frozen language model to predict. LLaMA-7B-ICL and Alpaca-7B-ICL using 2-shot samples with in-context learning to predict. Please refer to Section 4.1 for detailed introductions of these methods.
For structure information-based methods, we reproduce the results with OpenKE (Han et al., 2018). For additional information-based methods, we reproduce the results with official code implementation. For KGT5, there’s no public record or code for the method on the FB15K237N, WN18RR, UMLS dataset. Please refer to Appendix A.2 for detailed experimental settings.
Method Pred Samples FB13 WN11 15K237N WN18RR UMLS
TransE S-F ALL 0.815 0.759 0.697 0.884 0.845
DistMult S-F ALL 0.862 0.871 0.587 0.851 0.864
ComplEx S-F ALL 0.857 0.862 0.657 0.841 0.908
RotatE S-F ALL 0.819 0.847 0.685 0.882 0.921
KG-BERT L-MLP ALL 0.904 0.935 0.560 0.963 0.773
KGT5 N-L ALL 0.663 0.728 - - -
KG-LLAMA-7B N-L ALL 0.892 0.955 0.748 0.921 0.858
LLAMA-7B N-L 0 0.091 0.211 0.573 0.458 0.658
ALPACA-7B N-L 0 0.226 0.696 0.561 0.714 0.526
LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
ALPACA-7B-ICL N-L 2 0.675 0.706 0.607 0.691 0.605
ALPACA-DES N-L 0 0.611 - 0.590 0.548 0.525
LLAMA-MLP H-MLP 10,000 0.851 0.874 0.679 0.866 0.882
LLAMA-DES-H H-MLP 10,000 0.895 - 0.736 0.935 0.862
LLAMA-DES2 (Tri) H-MLP 10,000 0.847 0.764 0.673 0.785 0.782
LLAMA-DES2 (GPT)* H-MLP 1,000 0.890 0.892 0.705 0.907 0.783
Table 2: The main experiment results (accuracy) of triple classification. In the ’Pred’ column, S-F uses score function, L-MLP uses MLP on last layer hidden states, N-L uses natural language, and H-MLP uses MLP on intermediate layer hidden states to predict. ’Samples’ indicates training sample counts. ALPACA-DES combines non-generated descriptions with $PT_{2}$ for Alpaca-7B; LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimuli generation; LLAMA-DES-H combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-DES2 (Tri)/(GPT) uses first/second generated descriptions with $PT_{2}$ for stimuli generation. Top three results per dataset are bolded.
3.3 Main Results (For RQ1)
From Table 2, we can observe that LLAMA-DES-H and LLAMA-DES2 (GPT), when using a partial training set, exhibit comparable performance to KG-LLAMA-7B, the latter being fine-tuned with all training triples. Additionally, LLAMA-DES-H and LLAMA-DES2 (GPT) outperform LLAMA-MLP, demonstrating the effectiveness of using subgraph entity description generator to improve the LLM’s understanding of entities in the KG, as proposed in Section 2.3. Moreover, LLAMA-MLP demonstrates superior performance compared to LLAMA-7B, ALPACA-7B, LLAMA-7B-ICL and ALPACA-7B-ICL, affirming the effectiveness of stimulation method proposed in Section 2.2. However, it is noteworthy that ALPACA-DES fails to outperform ALPACA-7B, suggesting that mere incorporation of entity descriptions may not yield direct performance enhancements, considering the possibility of LLMs revealing the wrong responses and being influenced by hallucinations. Thus, the experimental results substantiate the effectiveness of our proposed methodology. Additionally, the frozen language model can also achieve similar KGC effects as fine-tuning the language model.
Notice that the performance on the UMLS dataset is special, where using generated descriptions impedes the classification accuracy of the model, and using the non-generated entity descriptions is also worse than not using any entity descriptions. It is postulated that the underlying reason may lie in the specialized nature of entity names in the UMLS dataset, resulting in minimal ambiguity. The entities in the knowledge graph have well-aligned with the internal entities of the LLM. Meanwhile, using entity descriptions may be affected by uncleaned description text, and descriptions generated by GPT-3.5-turbo may be influenced by hallucination, thereby affecting the classification effectiveness of the induced hidden states.
In addition to the triple classification, we also validate the effectiveness of the proposed probing method on the relation prediction task as shown in Table 3. The experimental results similarly demonstrate that the probing method can achieve the nearly same performance as full training set fine-tuning, even when using only 0.6% of the training data. Detailed methodology and analysis are provided in Appendix B.1.
Figure 3: Differences in KGC performance across different layers of the LLAMA-DES2 (GPT) [left] and Mistral-DES2 (GPT) [right]. The horizontal axis represents the position of the intermediate layers in LLaMA or Mistral (which have different backbones), and the vertical axis represents the dataset.
Method Pred Samples YAGO3-10
KG-BERT L-MLP ALL 0.6816
KGT5 N-L ALL 0.5714
ChatGLM-6B N-L 0 0.0658
KG-ChatGLM-6B N-L ALL 0.5662
LLaMA-7B N-L 0 0.0348
LLaMA-13B N-L 0 0.0040
KG-LLaMA-7B N-L 0 0.7028
LLAMA-MLP-7B H-MLP 6996 0.5968
LLAMA-DES2 (GPT)* H-MLP 6996 0.6824
Table 3: The experiment results (Hits@1) of relation prediction. LLAMA-MLP-7B refers to using LLaMA-7B as the base language model, using MLP as the hidden state classifier, and generating stimuli using $PT_{3}$; LLAMA-DES2 (GPT) refers to using the second generated description combined with $PT_{4}$ to generate stimuli.
Method FB13 WN11 FB15K-237N
LLAMA-SVM 0.824 0.869 0.599
LLAMA-LR 0.840 0.863 0.665
LLAMA-MLP 0.851 0.874 0.679
Table 4: The variation in KGC prediction performance across different datasets using SVM, Logistic Regression, MLP as classification models. All these methods use LLaMA-7B as the base model, and generate stimuli using $PT_{1}$.
Model Method FB13 WN11
LLaMA KG-LLAMA-7B 0.892 0.955
LLAMA-MLP 0.851 0.874
LLAMA-DES-H 0.895 -
LLAMA-DES2 (GPT)* 0.890 0.892
Mistral KG-MISTRAL-7B 0.891 0.962
MISTRAL-MLP 0.848 0.903
MISTRAL-DES-H 0.884 -
MISTRAL-DES2 (GPT)* 0.875 0.912
Gemma KG-GEMMA-7B 0.897 0.944
GEMMA-MLP 0.810 0.895
GEMMA-DES-H 0.895 -
GEMMA-DES2 (GPT)* 0.851 0.911
Table 5: The variation in KGC prediction performance across different datasets using LLaMA, Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024) as base models (which have different backbones). Please refer to Table 2 for the number of training samples used by different methods.
3.4 Ablation Study (For RQ2)
Through the analysis depicted in the left portion of Figure 3, it is observed that with an increase in the index of the intermediate layer, the efficacy of hidden states for triple classification demonstrates an initial incline followed by a subsequent decline. Intermediate layers closer to natural language text output generally exhibit better prediction performance compared to the lowest layers. This observation corroborates the notion posited by Geva et al. (2020) that language models encode knowledge differentially across their Feed-Forward layers, with higher layers assimilating information from lower layers. During pretraining, it is possible that lower layers may not have stored certain knowledge, hence they cannot generate hidden states with effective triple classification effects. Furthermore, the prediction performance of the layers closer to the natural language text output is generally worse than those in the middle layers. This phenomenon may be attributed to the hallucination issue inherent in LLaMA, wherein higher layers could potentially either omit or alter triple information, leading to a blurred distinction between positive and negative triple hidden states induced by stimuli. However, the lower layers of LLaMA can effectively gather the stimulus information required for the completion of the knowledge graph tasks. Akin observations are discernible on the right side of Figure 3 for the Mistral model, which has a different backbone from LLaMA. Meanwhile, as observed in Table 5, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning with only a small sample size.
Regarding classification models, a comparative analysis encompassing MLP, SVM, and Logistic Regression, as delineated in Table 4, is conducted. The result shows that MLP outperforms all three datasets compared to the other two classification models, while Logistic Regression also exhibits results close to MLP. This underscores the efficacy of our stimulus methodology, whereby hidden states of positive and negative samples within a high-dimensional space exhibit linear separability.
Figure 4: The LLAMA-DES-H prediction performance on KGC varies with the size of the training dataset. The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples, and the vertical axis represents the accuracy on the test set.
3.5 Efficiency Study (For RQ3)
From Figure 4, we can observe the following: For the FB13 dataset, using 400 training samples, which is 0.06% of the training set, achieves 98.3% of the performance of using the entire training set; For the FB15K-237N dataset, using 0.57% of the training set achieves 99.6% of the performance of using the entire training set; For the WN18RR dataset, using 0.46% of the training set achieves 98.8% of the performance of using the entire training set; For the UMLS dataset, using 9.58% of the dataset achieves 88.7% of the performance of using the entire training set. Our experimental results confirm that the language model is capable of generating effective hidden states through the techniques described in Sections 2.2 and 2.3, demonstrating its potential for data-efficient scenarios within the context of knowledge graph completion tasks.
Meanwhile, we also evaluate the compute efficiency of our method. Table 8 shows that our method can achieve significantly lower consumption compared to fine-tuning even in the setting of the full dataset. Compared to fine-tuned LLMs, we boost the GPU memory efficiency to 188×188$\times$188 × during model training and speed up 13.48×13.48$\times$13.48 × overall. Please refer to Appendix B.3 for details.
4 Related Work
4.1 Knowledge Graph Completion
Knowledge graphs are typically characterized by their inherent incompleteness, often containing substantial amounts of implicit or missing valuable knowledge (Hogan et al., 2021). The prevailing methodologies in KGC can be categorized into two main classes: structure information-based methods and additional information-based methods (Shen et al., 2022b). Structure information-based methods leverage the inherent structural data present within knowledge graphs. For example, RESCAL (Nickel et al., 2011), DistMult (Yang et al., 2014), COMPGCN (Vashishth et al., 2019), and ConvE (Dettmers et al., 2018) primarily use scoring functions to evaluate the semantic similarities of entity or relation embeddings within a latent embedding space to infer missing information. Additional information-based methods incorporate a variety of supplementary data to enrich the knowledge graphs. These methods utilize node attributes, entity-related information, or relational path data to enhance the features of knowledge embeddings critical for KGC tasks. The prominent examples are KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which assess the plausibility of triples by transforming the triple and its contextual data into natural language sentences using predefined templates. These sentences are then encoded using a fine-tuned BERT model, effectively leveraging additional information to improve the accuracy of triple classification. KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) employ encoder-decoder architecture language models to complete KGC in a generative manner. KG-LLAMA (Yao et al., 2023) utilizes triples composed of entity and relation names as sentences, fine-tuning the LLaMA model for KGC using natural language. Li et al. (2024) enhances KGC with LLM by generating entity descriptions and designing loss functions for fine-tuning.
4.2 Large Language Models
Several studies have explored the mechanisms through which Large Language Models process, interpret, and generate factual content. As highlighted by Saier et al. (2023), these models function as extensive repositories of knowledge, encapsulating a vast array of information about the world within their parameters. This insight into LLMs underscores their potential as powerful tools for a wide range of applications, including knowledge graph completion, by leveraging the rich informational content stored in their intricate network structures. Ramrakhiyani et al. (2023) have conducted evaluations across multiple Large Language Models to assess their knowledge of world geography using probing techniques. Similarly, Onoe et al. (2022) developed a methodology for automatically constructing cloze prompts by masking out spans, aiming to assess the language models’ understanding of entities through factual probing. In addition, Azaria and Mitchell (2023) and Zou et al. (2023) have explored training classifiers to estimate the truthfulness of statements based on the activations of the hidden layers in LLMs.
5 Conclusion
In this paper, we introduce a novel data-efficient probing method that completes KGC by leveraging the internal knowledge of LLMs. Experiments show that using the probing method with generated entity descriptions, frozen LLMs can achieve performance comparable to that of their fine-tuned counterparts. Simultaneously, this method saves a significant amount of GPU memory and time compared to fine-tuning approaches. This work explores the potential of frozen LLMs in KGC tasks with both efficacy and efficiency. Compared to fine-tuning LLMs, we hit the core of KGC problem in a better way.
Limitations
In this study, our method was primarily tested using the UMLS domain-specific knowledge graph and was not applied to other domain-specific KGs. Our current architectural design is specifically tailored for triple classification and relation prediction within knowledge graph completion. Recognizing these limitations, future work will aim to expand our architecture to accommodate additional KGC tasks such as link prediction. This extension will potentially broaden the applicability and enhance the robustness of our method across a more diverse set of knowledge graph scenarios.
Appendix A Experimental Details
A.1 Generate Entity Descriptions using GPT-3.5-turbo
In Section 2.3, we utilize LLM with the in-context learning technique to generate descriptions for entities based on the triples in their one-hop subgraph. In the experiments, we utilize the GPT-3.5 as the LLM. The prompt template used to generate entity descriptions is shown in Figure 5. Specifically, the $EN(e)$ is the entity name of entity $e$; the $\operatorname{concat\_subgraph()}$ is a function that first converts triples from the one-hop subgraph where the entity is located into sentences, and then concatenates these sentences using a specified delimiter. It is important to note that all the triples used in the prompt are from the training set, preventing potential data leakage from the test set in the generated entity descriptions. Table 6 shows the descriptions of entities generated using this method, thus achieving "entity alignment" between the knowledge graph and the LLM.
Entity Generated Description Non-Generated Description
Gustav Mahler
Gustav Mahler was a Jewish composer
known for his significant contributions
to the world of music. He lived in the
United States and was of German nationality.
Gustav Mahler (German; 7 July 1860 – 18 May 1911)
was an Austro-Bohemian late-Romantic composer,
and one of the leading conductors of his generation…
James Hayter
James Hayter was a male actor born in Scotland,
known for his profession in acting.
He passed away in Spain.
Henry James Goodenough Hayter
(23 April 1907 – 27 March 1983), better known
as James Hayter, was a British actor. He is best
remembered for his roles as Friar Tuck in the film The Story
of Robin Hood and His Merrie Men (1952) and…
Table 6: Comparison of generated entity descriptions with non-generated entity descriptions on the FB13 dataset.
Figure 5: The prompt template used to generate entity descriptions based on the one-hop subgraph where the entity is located for GPT-3.5-turbo.
A.2 Implementation Details
Figure 6: The prompt template used in Alpaca-7B-ICL on the WN18RR dataset.
In our implementation, we use the grid search to find appropriate sets of hyperparameters for baselines and our proposed method.
For the MLP classifier configuration, we opt for a batch size of 64, a learning rate of 3e-5, and utilize the AdamW optimizer, with training restricted to 30 epochs. For LLaMA-7B-ICL and Alpaca-7B-ICL, we randomly select one positive sample and one negative sample from the training set of each dataset, respectively. The prompt for KGC using Alpaca-7B-ICL on the WN18RR dataset is illustrated in Figure 6. Additionally, the hyper-parameters for models trained with the LoPA technique, specifically KG-LLAMA-7B, are outlined in Table 7.
It’s worth noting that for Table 2, Table 3, Table 4, Table 5 and Figure 4, we determine the most suitable layer for the task by experimenting with the validation set, and then report the results on the test set. For Figure 3, the results shown are for different layers of the language model on the test set. For the baseline results in Table 2 and Table 3, we give priority to using the results reported in the official baseline paper. For results not included in the paper, we replicate the experiments using the hyperparameters recommended by the official sources and conduct multiple attempts and optimizations on these hyperparameters as well. Each experiment is executed three times with random initializations, and the mean results are reported.
Parameter Value
optimizer adamw
learning rate 3e-4
epoch 3
weight decay 0
LoRA modules q_proj, k_proj, v_proj, o_proj
LoRA r 8
LoRA alpha 16
LoRA dropout 0.05
Table 7: Hyper-parameters of KG-LLAMA-7B.
Appendix B Additional Results and Analysis
B.1 Relation Prediction
Figure 7: Prompt templates to stimulate the LLM for relation prediction.
Method Procedure GPU Memory Time
KG-LLAMA
Training
(LLM:1 epoch)
14.68G
(LLM Parameters+LoRA+Gradient etc)
83h
(Forward+Backward)
Inference
(LLM:Generation)
12.94G
(LLM Parameters+LoRA)
2h50min
LLAMA-MLP
Training
(MLP:10 epoch)
0.078G
(MLP Parameters+Gradient etc)
33min
(Forward+Backward)
Inference
(LLM:Probing+MLP)
12.82G+0.078G
(LLM Parameters+MLP Parameters)
5h10min+29min+15s
Table 8: Efficiency Comparsion between KG-LLAMA and LLAMA-MLP on the WN11 dataset. The "Procedure" column represents several key steps of these methods.
Given an incomplete triple $(h,?,t)$ as query, relation prediction aims to predict the missing relation (denoted as ????). For relation prediction, we transform the binary framework of triple classification shown in Figure 1 into a multi-class framework. The categories for classification are the relationship types within each knowledge graph. Table 3 shows our experimental results on the YAGO3-10 (Dettmers et al., 2018) dataset. We use the following prompt template, $PT_{3}$ and $PT_{4}$, to create training samples for classification, as shown in Figure 7. $PT_{4}$ introduces the $D(e)$ to generate more precise and distinguishable hidden states, linking entities in the Yago with entities in the LLM, thereby improving the overall relation prediction performance.
It is worth noting that "ALL" represents 1,079,040 training samples. LLaMA-MLP and LLAMA-DES2 (GPT) only use 0.6% of the training set to train the probing classifier. It can be seen that LLAMA-DES2 (GPT) with frozen language model parameters achieves results very close to KG-LLAMA-7B, which is fine-tuned on the full dataset.
B.2 Case Study
Table 9 shows the prediction results of different models for two samples in the WN11 test set. For the first case, it can be seen that among the methods using frozen LLMs, only LLAMA-7B-ICL, LLAMA-MLP, and LLAMA-DES-H predict correctly. However, the LLAMA-7B-ICL method predicts "True" for almost all test samples. For the second case, it can be seen that only LLAMA-MLP, LLAMA-DES-H, and KG-LLAMA-7B predict correctly. However, KG-LLAMA-7B is fine-tuned on the full training set. The correct predictions by LLAMA-MLP and LLAMA-DES-H indicate that the LLM itself has the relevant knowledge to make accurate predictions, whereas methods like LLAMA-7B-ICL and ALPACA-7B do not fully utilize the knowledge stored in the language model or generate hallucinations during the response process.
QUERY MODEL PRED LABEL
HEAD: parsnip; REL: type of; TAIL: herb; LLAMA-7B N/A TRUE
LLAMA-7B-ICL TRUE
ALPACA-7B FALSE
ALPACA-7B-ICL FALSE
LLAMA-MLP TRUE
LLAMA-DES-H TRUE
KG-LLAMA-7B TRUE
HEAD: market synset; REL: domain topic; TAIL: brain stem; LLAMA-7B N/A FALSE
LLAMA-7B-ICL TRUE
ALPACA-7B TRUE
ALPACA-7B-ICL TRUE
LLAMA-MLP FALSE
LLAMA-DES-H FALSE
KG-LLAMA-7B FALSE
Table 9: Case study (triple classification) on two representative examples selected from the WN11 dataset. "N/A" indicates that the model’s output does not contain the predicted label information (i.e., yes or no).
B.3 Efficiency Analysis
We compared the performance of the KG-LLAMA method and LLAMA-MLP on the WN11 full dataset in terms of GPU memory and time consuming metrics using an Nvidia 3090 GPU. The results are shown in Table 8. For the computation of GPU memory usage, we use "torch.cuda.memory_allocated" to obtain the peak memory usage at each step. For the calculation of GPU memory reduction factor, we focus solely on the training phase, dividing the memory usage of KG-LLAMA by that of LLAMA-MLP. For the reduction factor of time consumption, we first sum the time consumed during both the training and inference phases for each method and then divide the results. For all methods, the batch size for all procedures is set to 1. The KG-LLAMA method employs LoRA for supervised fine-tuning, and its hyperparameters are listed in Table 7, with the epoch set to 1 in the efficiency experiment. For the training stage of KG-LLAMA, we employ 8-bit quantization technique, while for the inference stage, we configure not to use beam search. For the inference stage of LLAMA-MLP, we probe both the training data and the testing data. Thus, the sum of the three times in the table corresponds to the time required for probing the training data, probing the testing data, and the MLP forward pass, respectively.
It is worth noting that in practical use, our probing method only requires a small portion of the training set to complete the training of classifiers (such as MLP, Logistic Regression, etc.). Here, we used the full dataset for fairness in testing. We can clearly observe in the Table 8 that the proposed probing method, which uses frozen language models, consumes less GPU memory at all stages compared to KG-LLAMA, even though KG-LLAMA employs parameter-efficient fine-tuning techniques like LoRA to save memory. Additionally, since our method only requires a forward pass through the LLM, it consumes significantly less time compared to KG-LLAMA.
B.4 Analysis of Methods for Entity Description Generation
From Table 2, it can be observed that the performance of LLAMA-DES2 (Tri) is significantly inferior to that of LLAMA-DES2 (GPT), and even LLAMA-MLP. This indicates that using subgraph triples as entity descriptions directly does not enhance the model’s classification performance; it may even impair its judgment. This phenomenon may be attributed to subgraph triples diverging significantly from the language model’s pre-training corpus, thus providing erroneous stimulus information. However, organizing triples into coherent semantic natural language text using methods in Appendix A.1 appears to mitigate this impact.
B.5 Analysis of Using Frozen LLM in KGC
It has been observed that the predictive performance achieved by the frozen language models can not achieve comparable performance with supervised fine-tuning the language models (Zhang et al., 2023; Yao et al., 2023). But the knowledge graphs used in the experiments are encyclopedic knowledge graphs, such as FB15K-237 (Toutanova and Chen, 2015), which are subsets of FB15k (Bordes et al., 2013) derived from Wikipedia. Large Language Models like LLaMA (Touvron et al., 2023), which are pretrained on datasets such as Wikipedia, CommonCrawl, etc., have stored this type of world knowledge (Davison et al., 2019). Liu et al. (2023) shows that prompting can effectively access the knowledge within LLMs without fine-tuning. Therefore, frozen LLMs have sufficient knowledge and capability to complete KGC tasks.
Figure 8: We visualize the hidden states obtained from stimulation in the FB13 test set in three-dimensional space using PCA. The high dimensional vectors are obtained using the LLAMA-DES2 (GPT) method, collecting the language model’s hidden states from the 16th layer through stimulation.

### v2

@@
 Abstract
-Classical knowledge graph completion (KGC) methods rely solely on structural information, struggling with the inherent sparsity of knowledge graphs (KGs).
-Large Language Models (LLMs) learn extensive knowledge from large corpora with powerful context modeling, which is ideal for mitigating the limitations of previous methods.
+Traditional knowledge graph completion (KGC) methods rely solely on structural information, struggling with the inherent sparsity of knowledge graphs (KGs).
+Large Language Models (LLMs) learn extensive knowledge from large corpora with powerful context modeling, making them promising for mitigating the limitations of previous methods.
 Directly fine-tuning LLMs offers great capability but comes at the cost of huge time and memory consumption, while utilizing frozen LLMs yields suboptimal results.
@@
 We then train a data-efficient classifier on these hidden states to harness the inherent capabilities of frozen LLMs in KGC.
-We also generate entity descriptions with subgraph sampling on KGs, reducing the ambiguity of triplets and enriching the knowledge representation.
-Extensive experiments on standard benchmarks showcase the efficiency and effectiveness of our approach.
-We outperform classical KGC methods on most datasets and match the performance of fine-tuned LLMs.
-Additionally, compared to fine-tuned LLMs, we boost GPU memory efficiency by 188×188$\times$188 × and speed up training+inference by 13.48×13.48$\times$13.48 ×.
+Additionally, to reduce ambiguity and enrich knowledge representation, we generate detailed entity descriptions through subgraph sampling on KGs.
+Extensive experiments on standard benchmarks demonstrate the efficiency and effectiveness of our approach.
+We outperform traditional KGC methods across most datasets and, notably, achieve classification performance comparable to fine-tuned LLMs while enhancing GPU memory efficiency by 188×188$\times$188 × and accelerating training and inference by 13.48×13.48$\times$13.48 ×.
+Unlock the Power of Frozen LLMs in Knowledge Graph Completion
+Bo Xue, Yi Xu, Yunchong Song, Yiming Pang, Yuyang Ren, Jiaxin Dingthanks:  Jiaxin Ding is the corresponding authors., Luoyi Fu, Xinbing Wang Shanghai Jiao Tong University, Shanghai, China {sappho_x, jiaxinding, yiluofu} @sjtu.edu.cn
 1 Introduction
-Knowledge graph completion has become a crucial endeavor for the effective application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b).
-Classical knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which suffer from the sparseness of KGs (Yao et al., 2019).
+Knowledge graph completion has become a crucial endeavor to improve the application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b).
+Traditional knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which suffer from the sparseness of KGs (Yao et al., 2019).
 Leveraging the powerful contextual modeling capabilities and extensive stored world knowledge of LLMs can mitigate this challenge by utilizing the rich semantics within knowledge graphs, thereby enhancing the performance of knowledge graph completion (Wei et al., 2023).
-These approaches generally fall into two types: using frozen LLMs and employing fine-tuned LLMs.
-The first type leverages frozen language models (e.g., LLaMA and GPT-4) with prompt techniques, such as in-context learning, which facilitate the completion of knowledge graphs directly in natural language outputs (Wei et al., 2023; Yao et al., 2023), offering an efficient solution that utilizes the pre-trained capabilities of these models without the need for further training.
-The second type involves fine-tuning the language models using a range of supervised techniques, including instruction tuning, to tailor the models specifically for KGC tasks.
-This method involves fine-tuning language models, such as LLaMA and T5, through the use of training samples (Saxena et al., 2022; Yao et al., 2023).
-Although this can lead to better performance (Zhang et al., 2023; Yao et al., 2023), it also requires significant increases in time and resource expenditure.
-A natural question is: How can we utilize LLMs to accomplish KGC tasks with both efficacy and efficiency?
-To answer this question, we build upon the efficiency of frozen LLMs and explore the reasons behind the underperformance of frozen LLMs compared to their fine-tuned counterparts.
+These approaches generally fall into two types: using frozen LLMs with prompting techniques and fine-tuned LLMs specifically for KGC.
+The first type leverages frozen LLMs with prompt techniques, such as in-context learning.
+These techniques enable models to perform KGC directly through natural language outputs (Wei et al., 2023; Yao et al., 2023), harnessing the pre-trained knowledge embedded within these models without additional training.
+While this approach offers computational efficiency, it may not fully capture the task-specific nuances required for optimal KGC performance.
+The second type involves fine-tuning the LLMs using supervised techniques, including instruction tuning, to adapt models specifically for KGC tasks.
+This method involves fine-tuning LLMs, such as LLaMA, through the use of training samples (Yao et al., 2023), which can lead to improved performance by aligning the models more closely with the KGC objectives (Zhang et al., 2023; Yao et al., 2023).
+However, fine-tuning demands significant increases in time and resource expenditure.
+This dichotomy raises a fundamental question: How can we utilize LLMs to accomplish KGC tasks with both efficacy and efficiency?
+To answer this question, we explore the reasons behind the underperformance of frozen LLMs compared to their fine-tuned counterparts.
 First, while LLMs contain vast amounts of world knowledge encoded within their parameters, accessing the knowledge and decoding it into responses effectively through frozen models can be challenging (Wang et al., 2023).
-It is well known that LLMs suffer from hallucinations (Zhao et al., 2023), which makes them unreliable for achieving high performance in KGC tasks.
-Fine-tuning directly modifies the models’ weights through the training samples to improve the retrieval and application of relevant world knowledge, thus leading to better performance (Zhou et al., 2024).
-Second, KGs often involve complex relationships and ambiguous entity representations (Li et al., 2024), which can be difficult for a frozen LLM to handle accurately without specific tuning.
-Fine-tuned models can be explicitly trained to interpret and disambiguate such complexities on the basis of the training data provided, linking entities in the KG with the LLM.
-To overcome the limitations of frozen LLMs in KGC tasks, we have introduced probing techniques utilizing stimulation prompts, complemented by detailed descriptions of entities, as shown in Figure 1.
+LLMs can produce hallucinations-plausible but incorrect outputs (Zhao et al., 2023)-making them unreliable for precise KGC tasks.
+Fine-tuning directly modifies the models’ weights through task-specific data enhancing the retrieval of relevant knowledge and improving performance (Zhou et al., 2024).
+Second, KGs often involve complex relationships and ambiguous entity representations (Li et al., 2024), which can be difficult for a frozen LLM to handle such nuances accurately without specific tuning.
+Fine-tuned models can be explicitly trained to interpret and disambiguate the complexities enabling better linkage of entities within KGs.
+To overcome the limitations of frozen LLMs in KGC tasks, we introduce probing techniques that utilize stimulation prompts, complemented by detailed entity descriptions, as shown in Figure 1.
 These methods are designed to effectively extract and leverage the latent knowledge embedded within the models’ parameters, and minimize ambiguous entity representations, thereby enhancing its application in KGC without the need for extensive retraining.
-Specifically, our research concentrates on the core issues of knowledge graph completion, which are triple classification and relation prediction.
+Specifically, our research concentrates on the core tasks of knowledge graph completion: triple classification and relation prediction.
 We stimulate the intermediate layers of LLMs to capture the context-aware hidden states of knowledge triples.
 Moreover, we improve the precision of knowledge representation by aligning KG entities with the LLMs, reducing the ambiguity of knowledge triples through the generation of entity descriptions via subgraph sampling from KGs.
-We train a data-efficient classifier on these obtained hidden states suppervised by the labels, to fully harness the inherent KGC capabilities of the frozen LLMs.
+We train a data-efficient classifier on the obtained hidden states suppervised by labels, to fully harness the inherent KGC capabilities of the frozen LLMs.
 Figure 1: The overall architecture of our proposed probing method for the triple classification.
@@
 Our contributions can be summarized as follows.
-  • We introduce a novel probing method tailored for KGC tasks, designed to stimulate language models to verify knowledge graph triples directly.
+  • We introduce a novel probing method tailored for KGC tasks, designed to stimulate language models to predict knowledge graph triples directly.
 We further generate knowledge graph entity descriptions through the use of subgraphs, which are then seamlessly integrated into our probing techniques.
@@
 By utilizing only a subset of the dataset, our method achieves a relative performance improvement of up to 2.7% compared to fine-tuning the LLM on the entire training set.
-Additionally, it significantly reduces GPU memory usage by a factor of 188 during the training phase and descreases time consumption by a factor of 13.48, all while maintaining comparable performance.
+Additionally, it significantly reduces GPU memory usage by a factor of 188 during the training phase and decreases time consumption by a factor of 13.48, all while maintaining comparable performance.
 [^1] For a detailed comparison of efficiency, please refer to Table 8 in the Appendix B.3.
@@
 2 Method
-Our method’s design stems from an analysis of the two primary reasons for the underperformance of frozen LLMs compared to fine-tuned counterparts.
-First, LLM may not always trigger the depth of knowledge needed and reveal the right responses.
+Our method is designed based on an analysis of the two primary reasons for the underperformance of frozen LLMs compared to fine-tuned counterparts.
+First, frozen LLMs may not effectively access the necessary knowledge to generate accurate responses.
 We demonstrate in Section 2.2 how to achieve this for the frozen LLM using probing techniques.
@@
 2.1 Formulation and Notations
-A knowledge graph can be represented as a set of triples $G=\{(h,r,t)\}$, where $E$ and $R$ denote the set of entities and relations in G, respectively, $h\in E$ is a head entity, $t\in E$ is a tail entity, we collectively refer to the head entity and tail entity as $e$, and $r\in R$ represents the relation between them.
+A knowledge graph, $G$, comprises a set of entities, $E$, and relations, $R$.
+It can be represented as a set of triples $G=\{(h,r,t)\}$, where $h\in E$ is the head entity, $t\in E$ is the tail entity, we collectively refer to the head entity and tail entity as $e$, and $r\in R$ represents the relation between them.
 $D$ is the set of descriptions for each entity and relation.
@@
 Triple classification involves evaluating the plausibility of a given knowledge triple, to determine whether the triple is true or not and relation prediction predicts the missing relation in a triple.
-2.2 Stimulating the Intermediate Layers
+2.2 Stimulating the Intermediate Layers of Large Language Models
 Knowledge graphs consist of numerous triples, and the format of these triples is challenging for the LLM to understand, which may mislead the LLM to retrieve irrelevant knowledge and reveal the wrong responses.
-To minimize the influence of task-unrelated information and enhance the performance of LLMs in KGC, we leverage the internal judgment mechanisms of the language model.
-Our method involves strategically triggering the necessary depth of knowledge embedded within the LLM to elicit the correct responses required for accurate KGC.
-By focusing on extracting classification results directly from the model’s internal processing, we ensure that the responses are both relevant and precise, thereby improving the effectiveness of the KGC task.
-Since different layers of the language model have different roles (Azaria and Mitchell, 2023), we determine the most suitable layer for the task by experimenting with the validation set.
+To enhance the performance of LLMs in KGC, we leverage the internal judgment mechanisms of the language model (Zou et al., 2023).
+Our method involves strategically triggering the necessary depth of knowledge embedded within the LLM to elicit the correct responses required for accurate KGC, thereby improving the overall performance of the KGC task.
+Specifically, since different layers of the language model have different roles (Azaria and Mitchell, 2023), we determine the most suitable layer for the task by experimenting with the validation set.
+For the selection of intermediate layers, see Appendix B.5.
 First, we construct prompt templates to stimulate the model to generate the required classification information in the intermediate layers.
 In the triple classification, our goal is to determine whether the triples in the knowledge graph are true.
-Therefore, we use the following prompt template, $PT_{1}$, to create positive and negative sample pairs, as shown in Figure 2.
+Therefore, we use the following prompt template, $PT_{1}$, to create positive and negative sample pairs for training, as shown in Figure 2.
 Notice that the prompt does not describe positive or negative information for triples.
@@
 Specifically, we capture plausibility information for positive samples $s^{+}\in S^{+}$ by:
-$v\left(s^{+}\right)=\{M_{1-i}\left(s^{+}\right)[-1],0<i<L\},$ (1)
-where $M$ represents the language model, $M_{1-i}$ represents the first $i$ layers of the model, $[-1]$ refers to the hidden state corresponding to the last token, $L$ denotes the number of layers in the model.
-The dimension of $M_{1-i}(s^{+})[-1]$ is the embedding dimension of the language model.
+$v\left(s^{+}\right)=\{M_{1:i}\left(s^{+}\right)[-1],0<i<L\},$ (1)
+where $M$ represents the language model, $M_{1:i}$ represents the first $i$ layers of the model, $[-1]$ refers to the hidden state corresponding to the last token, $L$ denotes the number of layers in the model.
+The dimension of $M_{1:i}(s^{+})[-1]$ is the embedding dimension of the language model.
 Negative samples follow the same principle.
-At last, we train a classification model using $v(s^{+})$ and $v(s^{-})$ to capture plausibility information.
-The goal is to learn a boundary that can separate the hidden states of the intermediate layers of the language model corresponding to positive samples (which are true) from those corresponding to negative samples (which are false).
+At last, we train a classification model using $v(s^{+})$ and $v(s^{-})$ to distinguish between plausible and implausible triples.
+The goal is to learn a boundary that can separate the hidden states of the intermediate layers of the LLM corresponding to positive samples (which are true) from those corresponding to negative samples (which are false).
 We can use classification models such as multilayer perceptron (MLP), SVM, etc., for classification training.
@@
 where x represents $v(s)[l]$, $l$ is the layer index, $w$ and $b$ are weights and bias to be learned.
-For relation prediction, we simply transform the binary classification into multi-class classification.
+For relation prediction, we transform the binary classification into multi-class classification.
 By utilizing the prompt in Figure 7, we stimulate the LLM to generate hidden states that capture the characteristics of different relation types in the KG.
 Details can be found in the Appendix B.1.
-The task-unrelated information does not affect the classification results.
-The reason is that the positive samples $S^{+}$ and negative samples $S^{-}$ are constructed using the same prompt template.
-The classifier only needs to capture the plausibility of the triples in $v(S^{+})$ and $v(S^{-})$, that is, whether the triple is true or not.
-The advantage of using this method for KGC is that it can ignore the impact of the task-unrelated information and reveal the right responses for accurate KGC.
+By focusing on extracting classification results directly from the model’s internal processing, we ensure that the responses are both relevant and precise, thereby improving the effectiveness of the KGC task.
 Figure 2: Prompt templates to stimulate the LLM for triple classification.
 2.3 Subgraph Entity Description Generator
-The semantics of entities in the knowledge graph are highly complex; it is difficult for the language model to understand the entity based solely on the entity’s name.
+The semantics of entities in the knowledge graph are complex, making it challenging for the language model to interpret based solely on the entity’s name due to inherent ambiguity and polysemy.
 Therefore, we inform the language model about the meaning of each entity in the input.
@@
 We construct $D(e)$ in two ways.
-The first approach involves a straightforward method where we concatenate all the triples within the one-hop subgraph surrounding the entity $e$, thereby forming $D(e)$.
-However, this direct concatenation can sometimes lead to ineffective or misleading stimuli for the LLM, as the structure and content of these subgraph triples may significantly differ from the text seen during the LLM’s pre-training phase.
+The first approach involves a straightforward method where we concatenate all the triples within the one-hop subgraph surrounding the entity $e$, thereby forming $D(e)$ (Zhang et al., 2023).
+However, this direct concatenation can sometimes lead to ineffective or misleading stimuli for the LLM, as the structure of these subgraph triples may significantly differ from the text seen during the LLM’s pre-training phase.
 This divergence can adversely affect the model’s response accuracy.
@@
 where $\operatorname{subgraph(\cdot)}$ is a function that retrieves the set of all triples in the one-hop subgraph where the entity resides; $\operatorname{transform(\cdot)}$ is a function that converts the triple into sentence; $\operatorname{concat(\cdot)}$ is a function to process triple statements.
-For the first approach, $\operatorname{concat(\cdot)}$ is simply a string concatenation function that joins each transformed triple statement with a separator.
+For the first approach, $\operatorname{concat(\cdot)}$ is a string concatenation function that joins each transformed triple statement with a separator.
 For the second approach, we need to utilize LLM with the in-context learning technique (Brown et al., 2020) to generate descriptions of entities based on the triples in their subgraph.
@@
   • RQ2: How does our method perform across different language models and classification models?
-  • RQ3: How about data efficiency and compute efficiency of our proposed method?
+  • RQ3: What are the data efficiency and computation efficiency of our proposed method for KGC?
 3.1 Datasets
@@
 FB15K-237N 13,104 93 87,282 14,082 16,452
-WN18RR 40,943 11 86,835 3,034 3,134
+WN18RR 40,943 11 86,835 6,068 6,268
 UMLS 135 46 5,216 1,304 1,322
@@
 RotatE S-F ALL 0.819 0.847 0.685 0.882 0.921
-KG-BERT L-MLP ALL 0.904 0.935 0.560 0.963 0.773
+KG-BERT L-MLP ALL 0.904 0.935 0.560 0.927 0.773
 KGT5 N-L ALL 0.663 0.728 - - -
+LLAMA-7B N-L 0 0.699 0.661 0.573 0.458 0.658
+LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
 KG-LLAMA-7B N-L ALL 0.892 0.955 0.748 0.921 0.858
-LLAMA-7B N-L 0 0.091 0.211 0.573 0.458 0.658
-ALPACA-7B N-L 0 0.226 0.696 0.561 0.714 0.526
-LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
-ALPACA-7B-ICL N-L 2 0.675 0.706 0.607 0.691 0.605
-ALPACA-DES N-L 0 0.611 - 0.590 0.548 0.525
-LLAMA-MLP H-MLP 10,000 0.851 0.874 0.679 0.866 0.882
-LLAMA-DES-H H-MLP 10,000 0.895 - 0.736 0.935 0.862
-LLAMA-DES2 (Tri) H-MLP 10,000 0.847 0.764 0.673 0.785 0.782
-LLAMA-DES2 (GPT)* H-MLP 1,000 0.890 0.892 0.705 0.907 0.783
+KG-LLAMA-SAMPLED N-L 3k/10k 0.727 0.682 0.614 0.574 0.834
+LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
+LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.931 0.862
+LLAMA-DES2 (Tri) H-MLP 3k/10k 0.842 0.755 0.671 0.779 0.782
+LLAMA-DES2 (GPT)* H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
 Table 2: The main experiment results (accuracy) of triple classification.
-In the ’Pred’ column, S-F uses score function, L-MLP uses MLP on last layer hidden states, N-L uses natural language, and H-MLP uses MLP on intermediate layer hidden states to predict.
-’Samples’ indicates training sample counts.
-ALPACA-DES combines non-generated descriptions with $PT_{2}$ for Alpaca-7B; LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimuli generation; LLAMA-DES-H combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-DES2 (Tri)/(GPT) uses first/second generated descriptions with $PT_{2}$ for stimuli generation.
+In the $Pred$ column, $S\mbox{-}F$ uses score function, $L\mbox{-}MLP$ uses MLP on last layer hidden states, $N\mbox{-}L$ uses natural language, and $H\mbox{-}MLP$ uses MLP on intermediate layer hidden states to predict.
+$Samples$ indicates training sample counts, with $3k/10k$ indicating $10k$ samples for the UMLS dataset and $3k$ for others.
+LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimuli generation; LLAMA-DES-H combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-DES2 (Tri)/(GPT) uses first/second generated descriptions with $PT_{2}$ for stimuli generation.
 Top three results per dataset are bolded.
 3.3 Main Results (For RQ1)
-From Table 2, we can observe that LLAMA-DES-H and LLAMA-DES2 (GPT), when using a partial training set, exhibit comparable performance to KG-LLAMA-7B, the latter being fine-tuned with all training triples.
-Additionally, LLAMA-DES-H and LLAMA-DES2 (GPT) outperform LLAMA-MLP, demonstrating the effectiveness of using subgraph entity description generator to improve the LLM’s understanding of entities in the KG, as proposed in Section 2.3.
-Moreover, LLAMA-MLP demonstrates superior performance compared to LLAMA-7B, ALPACA-7B, LLAMA-7B-ICL and ALPACA-7B-ICL, affirming the effectiveness of stimulation method proposed in Section 2.2.
-However, it is noteworthy that ALPACA-DES fails to outperform ALPACA-7B, suggesting that mere incorporation of entity descriptions may not yield direct performance enhancements, considering the possibility of LLMs revealing the wrong responses and being influenced by hallucinations.
+From Table 2, it can be observed that LLAMA-DES-H and LLAMA-DES2 (GPT), when using a partial training set, exhibit comparable performance to KG-LLAMA-7B, the latter being fine-tuned with all training triples.
+Moreover, when trained on subsets of equivalent size, LLAMA-DES-H and LLAMA-DES2 (GPT) markedly surpass KG-LLAMA-SAMPLED, the existing state-of-the-art approach that relies exclusively on LLaMA for KGC.
+For more details, including the results of baseline models such as Alpaca-7B-ICL and Structural-aware-IT, and the performance of our method on the full training set, refer to Appendix B.4.
+The experimental results presented in Table 2 also validate the effectiveness of the approaches proposed in Section 2.2 and Section 2.3.
+Specifically, LLAMA-DES-H and LLAMA-DES2 (GPT) outperform LLAMA-MLP, demonstrating the effectiveness of using subgraph entity description generator to improve the LLM’s understanding of entities in the KG, as proposed in Section 2.3.
+Furthermore, LLAMA-MLP demonstrates superior performance compared to LLAMA-7B and LLAMA-7B-ICL, affirming the effectiveness of stimulation method proposed in Section 2.2.
 Thus, the experimental results substantiate the effectiveness of our proposed methodology.
-Additionally, the frozen language model can also achieve similar KGC effects as fine-tuning the language model.
-Notice that the performance on the UMLS dataset is special, where using generated descriptions impedes the classification accuracy of the model, and using the non-generated entity descriptions is also worse than not using any entity descriptions.
-It is postulated that the underlying reason may lie in the specialized nature of entity names in the UMLS dataset, resulting in minimal ambiguity.
+Additionally, the frozen LLM can also achieve similar KGC effects as that of fine-tuning the LLM.
+Notice that the performance on the UMLS dataset is special, where the use of generated descriptions hinders the model’s classification accuracy, and the use of non-generated entity descriptions is also worse than not using any entity descriptions.
+It is postulated that the underlying reason may lie in the highly specialized and domain-specific nature of entity names in the UMLS dataset, which are often precise medical terms with limited room for multiple interpretations.
 The entities in the knowledge graph have well-aligned with the internal entities of the LLM.
-Meanwhile, using entity descriptions may be affected by uncleaned description text, and descriptions generated by GPT-3.5-turbo may be influenced by hallucination, thereby affecting the classification effectiveness of the induced hidden states.
+Meanwhile, using entity descriptions may be affected by noisy description text, and descriptions generated by GPT-3.5-turbo may be influenced by hallucination, thereby affecting the classification effectiveness of the induced hidden states.
 In addition to the triple classification, we also validate the effectiveness of the proposed probing method on the relation prediction task as shown in Table 3.
@@
 LLAMA-MLP-7B refers to using LLaMA-7B as the base language model, using MLP as the hidden state classifier, and generating stimuli using $PT_{3}$; LLAMA-DES2 (GPT) refers to using the second generated description combined with $PT_{4}$ to generate stimuli.
+3.4 Ablation Study (For RQ2)
+Through the analysis depicted in the left portion of Figure 3, it is observed that with an increase in the index of the intermediate layer, the efficacy of hidden states for triple classification demonstrates an initial incline followed by a subsequent decline.
+Intermediate layers closer to natural language text output generally exhibit better prediction performance compared to the lowest layers.
+This observation corroborates the notion posited by Geva et al.
+(2020) that language models encode knowledge differentially across their Feed-Forward layers, with higher layers integrating information from lower layers.
+During pretraining, the lower layers may not have stored certain knowledge, and thus are unable to generate hidden states conducive to accurate triple classification.
+Furthermore, the prediction performance of the layers closer to the natural language text output is generally worse than those in the middle layers.
+This phenomenon may be attributed to the hallucination issue inherent in LLaMA, wherein higher layers could potentially either omit or alter triple information, leading to a blurred distinction between positive and negative triple hidden states induced by stimuli (Jiang et al., 2024; Zou et al., 2023).
+Detailed explanations can be found in Appendix B.5.
+Akin observations are discernible on the right side of Figure 3 for the Mistral model, which has a different backbone from LLaMA.
+Meanwhile, as observed in Table 5, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning with only a small sample size.
+Regarding classification models, a comparative analysis encompassing MLP, SVM, and Logistic Regression, as delineated in Table 4, is conducted.
+The result shows that MLP outperforms all three datasets compared to the other two classification models, while Logistic Regression also exhibits results close to MLP.
+This underscores the efficacy of our stimulus methodology, whereby hidden states of positive and negative samples within a high-dimensional space are easily distinguishable.
 Method FB13 WN11 FB15K-237N
-LLAMA-SVM 0.824 0.869 0.599
-LLAMA-LR 0.840 0.863 0.665
-LLAMA-MLP 0.851 0.874 0.679
+LLAMA-SVM 0.821 0.866 0.594
+LLAMA-LR 0.839 0.860 0.662
+LLAMA-MLP 0.855 0.872 0.674
 Table 4: The variation in KGC prediction performance across different datasets using SVM, Logistic Regression, MLP as classification models.
@@
 LLaMA KG-LLAMA-7B 0.892 0.955
-LLAMA-MLP 0.851 0.874
-LLAMA-DES-H 0.895 -
-LLAMA-DES2 (GPT)* 0.890 0.892
+LLAMA-MLP 0.855 0.872
+LLAMA-DES-H 0.901 -
+LLAMA-DES2 (GPT)* 0.912 0.917
 Mistral KG-MISTRAL-7B 0.891 0.962
-MISTRAL-MLP 0.848 0.903
-MISTRAL-DES-H 0.884 -
-MISTRAL-DES2 (GPT)* 0.875 0.912
+MISTRAL-MLP 0.846 0.901
+MISTRAL-DES-H 0.880 -
+MISTRAL-DES2 (GPT)* 0.893 0.925
 Gemma KG-GEMMA-7B 0.897 0.944
-GEMMA-MLP 0.810 0.895
-GEMMA-DES-H 0.895 -
-GEMMA-DES2 (GPT)* 0.851 0.911
+GEMMA-MLP 0.806 0.893
+GEMMA-DES-H 0.891 -
+GEMMA-DES2 (GPT)* 0.867 0.917
 Table 5: The variation in KGC prediction performance across different datasets using LLaMA, Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024) as base models (which have different backbones).
 Please refer to Table 2 for the number of training samples used by different methods.
-3.4 Ablation Study (For RQ2)
-Through the analysis depicted in the left portion of Figure 3, it is observed that with an increase in the index of the intermediate layer, the efficacy of hidden states for triple classification demonstrates an initial incline followed by a subsequent decline.
-Intermediate layers closer to natural language text output generally exhibit better prediction performance compared to the lowest layers.
-This observation corroborates the notion posited by Geva et al.
-(2020) that language models encode knowledge differentially across their Feed-Forward layers, with higher layers assimilating information from lower layers.
-During pretraining, it is possible that lower layers may not have stored certain knowledge, hence they cannot generate hidden states with effective triple classification effects.
-Furthermore, the prediction performance of the layers closer to the natural language text output is generally worse than those in the middle layers.
-This phenomenon may be attributed to the hallucination issue inherent in LLaMA, wherein higher layers could potentially either omit or alter triple information, leading to a blurred distinction between positive and negative triple hidden states induced by stimuli.
-However, the lower layers of LLaMA can effectively gather the stimulus information required for the completion of the knowledge graph tasks.
-Akin observations are discernible on the right side of Figure 3 for the Mistral model, which has a different backbone from LLaMA.
-Meanwhile, as observed in Table 5, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning with only a small sample size.
-Regarding classification models, a comparative analysis encompassing MLP, SVM, and Logistic Regression, as delineated in Table 4, is conducted.
-The result shows that MLP outperforms all three datasets compared to the other two classification models, while Logistic Regression also exhibits results close to MLP.
-This underscores the efficacy of our stimulus methodology, whereby hidden states of positive and negative samples within a high-dimensional space exhibit linear separability.
-Figure 4: The LLAMA-DES-H prediction performance on KGC varies with the size of the training dataset.
-The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples, and the vertical axis represents the accuracy on the test set.
 3.5 Efficiency Study (For RQ3)
-From Figure 4, we can observe the following: For the FB13 dataset, using 400 training samples, which is 0.06% of the training set, achieves 98.3% of the performance of using the entire training set; For the FB15K-237N dataset, using 0.57% of the training set achieves 99.6% of the performance of using the entire training set; For the WN18RR dataset, using 0.46% of the training set achieves 98.8% of the performance of using the entire training set; For the UMLS dataset, using 9.58% of the dataset achieves 88.7% of the performance of using the entire training set.
-Our experimental results confirm that the language model is capable of generating effective hidden states through the techniques described in Sections 2.2 and 2.3, demonstrating its potential for data-efficient scenarios within the context of knowledge graph completion tasks.
+As illustrated in Figure 4, our method notably surpasses both KG-LLAMA and KG-BERT on the FB13 when trained with only 100 samples, as opposed to the 3000 samples used by the latter methods.
+For the UMLS, the performance of KG-LLAMA and KG-BERT with 3000 training samples is comparable to that of LLAMA-DES-H trained with only 500 samples.
+Additionally, combining the results from Table 2 and Table 10, for the FB13 dataset, our method attains 97.2% of the performance achievable with the full training set, utilizing only 500 samples, which constitutes a mere 0.079% of the total training samples.
+In contrast, KG-LLAMA and KG-BERT achieve only 78.7% and 63.1% of the maximum performance, respectively.
+For the UMLS, with 9.58% of the training set, our method reaches 90.6% of the full training set performance.
+For other datasets, please refer to Appendix B.6.
+Our experimental results confirm that the language model is capable of generating effective hidden states through the techniques described in Sections 2.2 and 2.3, demonstrating its potential for data-efficient scenarios within the context of KGC tasks.
+Figure 4: The prediction performance of KG-LLAMA, KG-BERT, and LLAMA-DES-H on triple classification varies with the size of the training dataset.
+The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples.
 Meanwhile, we also evaluate the compute efficiency of our method.
-Table 8 shows that our method can achieve significantly lower consumption compared to fine-tuning even in the setting of the full dataset.
+Table 8 shows that our method can achieve significantly lower consumption compared to parameter-efficient fine-tuning even in the setting of the full dataset.
 Compared to fine-tuned LLMs, we boost the GPU memory efficiency to 188×188$\times$188 × during model training and speed up 13.48×13.48$\times$13.48 × overall.
@@
 These methods utilize node attributes, entity-related information, or relational path data to enhance the features of knowledge embeddings critical for KGC tasks.
-The prominent examples are KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which assess the plausibility of triples by transforming the triple and its contextual data into natural language sentences using predefined templates.
-These sentences are then encoded using a fine-tuned BERT model, effectively leveraging additional information to improve the accuracy of triple classification.
+The prominent examples are KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which represent entities, relations, and triples as textual sequences, turning KGC into a sequence classification problem.
 KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) employ encoder-decoder architecture language models to complete KGC in a generative manner.
-KG-LLAMA (Yao et al., 2023) utilizes triples composed of entity and relation names as sentences, fine-tuning the LLaMA model for KGC using natural language.
+KG-LLAMA (Yao et al., 2023) utilizes triples composed of entity and relation names as sentences, parameter-efficient fine-tuning the LLaMA model for KGC.
 Li et al.
@@
 Experiments show that using the probing method with generated entity descriptions, frozen LLMs can achieve performance comparable to that of their fine-tuned counterparts.
-Simultaneously, this method saves a significant amount of GPU memory and time compared to fine-tuning approaches.
+Simultaneously, this method saves a significant amount of GPU memory and time compared to parameter-efficient fine-tuning approaches.
 This work explores the potential of frozen LLMs in KGC tasks with both efficacy and efficiency.
@@
 Limitations
-In this study, our method was primarily tested using the UMLS domain-specific knowledge graph and was not applied to other domain-specific KGs.
-Our current architectural design is specifically tailored for triple classification and relation prediction within knowledge graph completion.
-Recognizing these limitations, future work will aim to expand our architecture to accommodate additional KGC tasks such as link prediction.
-This extension will potentially broaden the applicability and enhance the robustness of our method across a more diverse set of knowledge graph scenarios.
+Our current architectural design is only applicable to open-source LLMs or closed-source models that provide access to token embeddings; it is not suitable for closed-source models that exclusively output text.
+Recognizing these limitations, future work will aim to expand our architecture to accommodate closed-source LLMs that only output text or token probabilities.
+Additionally, future efforts could integrate structure information-based KG models, such as TransE, with our methodology in a hybrid approach.
 Appendix A Experimental Details
@@
 Figure 6: The prompt template used in Alpaca-7B-ICL on the WN18RR dataset.
-In our implementation, we use the grid search to find appropriate sets of hyperparameters for baselines and our proposed method.
+In our implementation, we experimented with these baselines using the hyperparameters recommended in their paper.
+For the negative sampling employed on the datasets, we utilize the same negative sampling technique as Yao et al.
+(2019) for FB13 and WN11.
+Specifically, FB15K-237N is designed to extract hard negative triples, providing a more challenging evaluation and reducing false negatives in the validation/test datasets during dataset construction.
+A similar approach is adopted for constructing negative samples for WN18RR and UMLS.
+Meanwhile, for all experiments in our study, the training sets contain an equal number of positive and negative samples.
 For the MLP classifier configuration, we opt for a batch size of 64, a learning rate of 3e-5, and utilize the AdamW optimizer, with training restricted to 30 epochs.
@@
 The prompt for KGC using Alpaca-7B-ICL on the WN18RR dataset is illustrated in Figure 6.
-Additionally, the hyper-parameters for models trained with the LoPA technique, specifically KG-LLAMA-7B, are outlined in Table 7.
+Additionally, the hyper-parameters for models trained with the LoRA technique, specifically KG-LLAMA-7B, are outlined in Table 7.
 It’s worth noting that for Table 2, Table 3, Table 4, Table 5 and Figure 4, we determine the most suitable layer for the task by experimenting with the validation set, and then report the results on the test set.
@@
 Additionally, since our method only requires a forward pass through the LLM, it consumes significantly less time compared to KG-LLAMA.
-B.4 Analysis of Methods for Entity Description Generation
+Method Pred Samples FB13 WN11 15K237N WN18RR UMLS
+LLAMA-7B N-L 0 0.699 0.661 0.573 0.458 0.658
+ALPACA-7B N-L 0 0.226 0.696 0.561 0.714 0.526
+LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
+ALPACA-7B-ICL N-L 2 0.675 0.706 0.607 0.691 0.605
+STRUCTURAL-AWARE-IT N-L ALL 0.907 0.959 0.752 0.927 0.881
+ALPACA-DES N-L 0 0.611 - 0.590 0.548 0.525
+LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
+LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.931 0.862
+LLAMA-DES2 (GPT)* H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
+LLAMA-DES-H H-MLP ALL 0.905 - 0.742 0.937 0.866
+Table 10: The experiment results (accuracy) of triple classification. ALPACA-DES combines non-generated descriptions with $PT_{2}$ for Alpaca-7B.
+B.4 Additional Baselines
+From Table 10, it is noteworthy that ALPACA-DES fails to outperform ALPACA-7B, suggesting that mere incorporation of entity descriptions may not yield direct performance enhancements, considering the possibility of LLMs revealing the wrong responses and being influenced by hallucinations.
+Meanwhile, in our study, we did not compare our method with KOPA because it integrates KG models (such as TransE/RotatE) with language models (like LLaMA) in a hybrid approach, thereby blending multiple approaches rather than solely utilizing the capabilities of LLMs (Zhang et al., 2023).
+In contrast, our current research exclusively focuses on the utilization of language models (like LLaMA), rendering KOPA an unsuitable baseline for our purposes.
+Nonetheless, we reproduced the Structural-aware-IT model proposed by Zhang et al.
+(2023), which incorporates in-context learning and instruction tuning.
+This model samples the neighborhoods of the head entity $h$ and tail entity $t$, including the textual descriptions of neighboring triples in the demonstration prompt, followed by fine-tuning on the entire training set.
+Like KOPA, this approach integrates structural information from the KG with LLM capabilities.
+Despite this, our method, which exclusively utilizes LLMs without integrating KG structural information, achieves performance comparable to the Structural-aware-IT model, which combines both KG structure and LLM capabilities.
+B.5 Analysis of the Performance in the Intermediate Layers of LLM
+The selection of the optimal layer index in LLMs serves as a hyperparameter, which we may select based on empirical evidence or adjust as needed when a validation set is available.
+Meanwhile, our experiments found a strong consistency for the same LLM across different knowledge graphs (for example, for LLAMA, the 15th layer often yields the best KGC performance), as shown in Figure 3, which reduces the complexity of utilization.
+Additionally, the superior performance of the intermediate layers of language models, as compared to those closer to the output layer, may be attributed to the potential for each layer of the model to experience hallucinations.
+Moreover, each layer of a language model can potentially propagate hallucinations from previous layers, manifesting as deliberate falsehoods or the expression of misleading information (Zou et al., 2023).
+This phenomenon can lead to scenarios where models often know the correct answer even while generating incorrect outputs (Zou et al., 2023).
+Concurrently, our experimental results are consistent with the conclusions of Jiang et al.
+(2024), which found that the MLP layers in the later stages have an inhibitory effect on the extracted knowledge, particularly contributing significantly to erroneous output decoding.
+As a result, the middle layers in their experiments also tend to outperform the later layers.
+B.6 Additional Analysis of the Data Efficiency
+From Figure 8, we can observe the following: For the FB13 dataset, using 400 training samples, which is 0.06% of the training set, achieves 97.0% of the performance of using the entire training set; For the FB15K-237N dataset, using 0.57% of the training set achieves 99.6% of the performance of using the entire training set; For the WN18RR dataset, using 0.46% of the training set achieves 98.8% of the performance of using the entire training set; For the UMLS dataset, using 9.58% of the dataset achieves 90.6% of the performance of using the entire training set.
+Figure 8: The LLAMA-DES-H prediction performance on KGC varies with the size of the training dataset.
+The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples, and the vertical axis represents the accuracy on the test set.
+It can be observed that the LLAMA-DES-H method exhibits slightly lower data efficiency on the UMLS dataset compared to other methods.
+This phenomenon may be attributed to the fact that UMLS is a domain-specific knowledge graph, whereas the other datasets are encyclopedic knowledge graphs.
+The increased complexity of the decision boundary in the hidden states of the positive and negative example triplets generated by LLaMA necessitates a larger volume of data to adequately model the boundary.
+B.7 Analysis of Methods for Entity Description Generation
 From Table 2, it can be observed that the performance of LLAMA-DES2 (Tri) is significantly inferior to that of LLAMA-DES2 (GPT), and even LLAMA-MLP.
@@
 However, organizing triples into coherent semantic natural language text using methods in Appendix A.1 appears to mitigate this impact.
-B.5 Analysis of Using Frozen LLM in KGC
+Figure 9: We visualize the hidden states obtained from stimulation in the FB13 test set in three-dimensional space using PCA.
+The high dimensional vectors are obtained using the LLAMA-DES2 (GPT) method, collecting the language model’s hidden states from the 16th layer through stimulation.
+B.8 Analysis of Using Frozen LLM in KGC
 It has been observed that the predictive performance achieved by the frozen language models can not achieve comparable performance with supervised fine-tuning the language models (Zhang et al., 2023; Yao et al., 2023).
 But the knowledge graphs used in the experiments are encyclopedic knowledge graphs, such as FB15K-237 (Toutanova and Chen, 2015), which are subsets of FB15k (Bordes et al., 2013) derived from Wikipedia.
-Large Language Models like LLaMA (Touvron et al., 2023), which are pretrained on datasets such as Wikipedia, CommonCrawl, etc., have stored this type of world knowledge (Davison et al., 2019).
+LLMs like LLaMA (Touvron et al., 2023), which are pretrained on datasets such as Wikipedia, CommonCrawl, etc., have stored this type of world knowledge (Davison et al., 2019).
 Liu et al.
@@
 Therefore, frozen LLMs have sufficient knowledge and capability to complete KGC tasks.
-Figure 8: We visualize the hidden states obtained from stimulation in the FB13 test set in three-dimensional space using PCA.
-The high dimensional vectors are obtained using the LLAMA-DES2 (GPT) method, collecting the language model’s hidden states from the 16th layer through stimulation.

### v3

@@
-Unlock the Power of Frozen LLMs in Knowledge Graph Completion
-Abstract
+Bridging LLMs and KGs without Fine-Tuning: Intermediate Probing Meets Subgraph-Aware Entity Descriptions
+Abstract.
 Traditional knowledge graph completion (KGC) methods rely solely on structural information, struggling with the inherent sparsity of knowledge graphs (KGs).
-Large Language Models (LLMs) learn extensive knowledge from large corpora with powerful context modeling, making them promising for mitigating the limitations of previous methods.
-Directly fine-tuning LLMs offers great capability but comes at the cost of huge time and memory consumption, while utilizing frozen LLMs yields suboptimal results.
-In this work, we aim to leverage LLMs for KGC effectively and efficiently.
-We capture the context-aware hidden states of knowledge triples by employing prompts to stimulate the intermediate layers of LLMs.
-We then train a data-efficient classifier on these hidden states to harness the inherent capabilities of frozen LLMs in KGC.
-Additionally, to reduce ambiguity and enrich knowledge representation, we generate detailed entity descriptions through subgraph sampling on KGs.
-Extensive experiments on standard benchmarks demonstrate the efficiency and effectiveness of our approach.
-We outperform traditional KGC methods across most datasets and, notably, achieve classification performance comparable to fine-tuned LLMs while enhancing GPU memory efficiency by 188×188$\times$188 × and accelerating training and inference by 13.48×13.48$\times$13.48 ×.
-Unlock the Power of Frozen LLMs in Knowledge Graph Completion
-Bo Xue, Yi Xu, Yunchong Song, Yiming Pang, Yuyang Ren, Jiaxin Dingthanks:  Jiaxin Ding is the corresponding authors., Luoyi Fu, Xinbing Wang Shanghai Jiao Tong University, Shanghai, China {sappho_x, jiaxinding, yiluofu} @sjtu.edu.cn
-1 Introduction
+By contrast, Large Language Models (LLMs) encapsulate extensive world knowledge and exhibit powerful context modeling capabilities, making them promising for mitigating the limitations of traditional methods.
+However, direct fine-tuning of LLMs for KGC, though effective, imposes substantial computational and memory overheads, while utilizing non-fine-tuned LLMs is efficient but yields suboptimal performance.
+In this work, we propose a novel framework that synergizes the strength of LLMs with robust knowledge representation to enable effective and efficient KGC.
+We extract the context-aware hidden states of knowledge triples from the intermediate layers of LLMs, thereby capturing rich semantic and relational nuances.
+These represenation are then utilized to train a data-effecient classifier tailored specifically for KGC tasks.
+To bridge the semanitic gaps between LLMs and KGs, we employ subgraph sampling on KGs to generate model-friendly entity descriptions, ensuring that the representations are comprehensive and well aligned with the KG structure.
+Extensive experiments on standard benchmarks validate the efficiency and effectiveness of our approach.
+We achieve a 47% relative improvement over previous methods based on non-fine-tuned LLMs and, to our knowledge, are the first to achieve classification performance comparable to fine-tuned LLMs while enhancing GPU memory efficiency by 188×188$\times$188 × and accelerating training and inference by 26.11×26.$1$.
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
+1. Introduction
 Knowledge graph completion has become a crucial endeavor to improve the application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b).
-Traditional knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which suffer from the sparseness of KGs (Yao et al., 2019).
-Leveraging the powerful contextual modeling capabilities and extensive stored world knowledge of LLMs can mitigate this challenge by utilizing the rich semantics within knowledge graphs, thereby enhancing the performance of knowledge graph completion (Wei et al., 2023).
-These approaches generally fall into two types: using frozen LLMs with prompting techniques and fine-tuned LLMs specifically for KGC.
-The first type leverages frozen LLMs with prompt techniques, such as in-context learning.
-These techniques enable models to perform KGC directly through natural language outputs (Wei et al., 2023; Yao et al., 2023), harnessing the pre-trained knowledge embedded within these models without additional training.
+Traditional knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which is vulnerable to the sparseness of KGs (Yao et al., 2019).
+By leveraging the powerful contextual modeling capabilities and extensive stored world knowledge of LLMs, we can effectively mitigate these challenges, and enhance the performance of KGC (Wei et al., 2024).
+These approaches generally fall into two types: using non-fine-tuned LLMs with prompting techniques and fine-tuned LLMs specifically for KGC.
+The first type leverages LLMs with prompt techniques, such as in-context learning.
+These techniques enable models to perform KGC directly through natural language outputs (Yang et al., 2024; Wei et al., 2024; Yao et al., 2024), harnessing the pre-trained knowledge embedded within these models without additional training.
 While this approach offers computational efficiency, it may not fully capture the task-specific nuances required for optimal KGC performance.
-The second type involves fine-tuning the LLMs using supervised techniques, including instruction tuning, to adapt models specifically for KGC tasks.
-This method involves fine-tuning LLMs, such as LLaMA, through the use of training samples (Yao et al., 2023), which can lead to improved performance by aligning the models more closely with the KGC objectives (Zhang et al., 2023; Yao et al., 2023).
+In contrast, the latter type involves fine-tuning the LLMs using supervised techniques to adapt models specifically for KGC tasks.
+This method involves fine-tuning LLMs, such as LLaMA, through the use of training samples (Yao et al., 2024), which can lead to improved performance by aligning the models more closely with the KGC objectives (Zhang et al., 2024; Yao et al., 2024).
 However, fine-tuning demands significant increases in time and resource expenditure.
-This dichotomy raises a fundamental question: How can we utilize LLMs to accomplish KGC tasks with both efficacy and efficiency?
-To answer this question, we explore the reasons behind the underperformance of frozen LLMs compared to their fine-tuned counterparts.
-First, while LLMs contain vast amounts of world knowledge encoded within their parameters, accessing the knowledge and decoding it into responses effectively through frozen models can be challenging (Wang et al., 2023).
-LLMs can produce hallucinations-plausible but incorrect outputs (Zhao et al., 2023)-making them unreliable for precise KGC tasks.
-Fine-tuning directly modifies the models’ weights through task-specific data enhancing the retrieval of relevant knowledge and improving performance (Zhou et al., 2024).
-Second, KGs often involve complex relationships and ambiguous entity representations (Li et al., 2024), which can be difficult for a frozen LLM to handle such nuances accurately without specific tuning.
-Fine-tuned models can be explicitly trained to interpret and disambiguate the complexities enabling better linkage of entities within KGs.
-To overcome the limitations of frozen LLMs in KGC tasks, we introduce probing techniques that utilize stimulation prompts, complemented by detailed entity descriptions, as shown in Figure 1.
-These methods are designed to effectively extract and leverage the latent knowledge embedded within the models’ parameters, and minimize ambiguous entity representations, thereby enhancing its application in KGC without the need for extensive retraining.
-Specifically, our research concentrates on the core tasks of knowledge graph completion: triple classification and relation prediction.
-We stimulate the intermediate layers of LLMs to capture the context-aware hidden states of knowledge triples.
-Moreover, we improve the precision of knowledge representation by aligning KG entities with the LLMs, reducing the ambiguity of knowledge triples through the generation of entity descriptions via subgraph sampling from KGs.
-We train a data-efficient classifier on the obtained hidden states suppervised by labels, to fully harness the inherent KGC capabilities of the frozen LLMs.
-Figure 1: The overall architecture of our proposed probing method for the triple classification.
-Constructing positive and negative sample pairs, followed by stimulating the middle layers of the language model to obtain hidden states for completing the KGC task.
-Entity descriptions are generated by the subgraph entity description generator.
-The only component that needs training in the entire architecture is the data-efficient classifier used to classify hidden states, which can be MLP, logistic regression, etc.
-Our contributions can be summarized as follows.
-  • We introduce a novel probing method tailored for KGC tasks, designed to stimulate language models to predict knowledge graph triples directly.
-We further generate knowledge graph entity descriptions through the use of subgraphs, which are then seamlessly integrated into our probing techniques.
-  • Our approach employs LLMs with both effectiveness and efficiency to achieve KGC.
-By utilizing only a subset of the dataset, our method achieves a relative performance improvement of up to 2.7% compared to fine-tuning the LLM on the entire training set.
-Additionally, it significantly reduces GPU memory usage by a factor of 188 during the training phase and decreases time consumption by a factor of 13.48, all while maintaining comparable performance.
-[^1] For a detailed comparison of efficiency, please refer to Table 8 in the Appendix B.3.
-  • We conduct extensive experiments on six standard knowledge graph datasets to benchmark our method against multiple baseline approaches.
-We investigate the versatility of our method with various language models, assessing its efficacy and data efficiency through rigorous experimental validation.
-2 Method
-Our method is designed based on an analysis of the two primary reasons for the underperformance of frozen LLMs compared to fine-tuned counterparts.
-First, frozen LLMs may not effectively access the necessary knowledge to generate accurate responses.
-We demonstrate in Section 2.2 how to achieve this for the frozen LLM using probing techniques.
-Second, KGs often involve ambiguous entity representations, which can be difficult for LLM to handle accurately.
-In Section 2.3 we show how to achieve this with the generation of entity descriptions.
-2.1 Formulation and Notations
+This dichotomy between computational efficiency and task-specific performance raises a fundamental question: How can we harness LLMs for effective and efficient KGC without incurring the high costs associated with fine-tuning?
+Answering this question necessitates a deeper look at the limitations of LLMs compared to their fine-tuned counterparts.
+Non-fine-tuned LLMs, while cost-effective, risk generating misleading outputs in the form of hallucinations (Zhao et al., 2023) and underutilizing their latent knowledge (Wang et al., 2023).
+Moreover, the intricate relational structures and semantic subtleties embedded in KGs compound the challenge, since LLMs are not inherently designed to parse or disambiguate complex entity linkages, without explicitly fine-tuning on such tasks (Li et al., 2024).
+In this paper, we propose a novel framework that reconciles these competing demands by fully exploiting the knowledge representation capabilities of LLMs without fine-tuning, as shown in Figure 1.
+Our method strategically extracts intermediate hidden states from LLMs to capture rich, context-aware representations of knowledge triples, which are then used to train a lightweight, data-efficient classifier specifically tailored for KGC.
+To bridge the gap between LLMs and KGs, we incorporate an innovative subgraph sampling technique to generate model-friendly entity descriptions, enabling the LLMs to better interpret and disambiguate relational patterns to further enhance knowledge representation.
+By focusing on these intermediate LLM representations, our framework mitigates hallucination risks while retaining the benefits of a non-fine-tuned paradigm, allowing for both high accuracy and reduced computational costs.
+Extensive experiments demonstrate that our framework not only outperforms existing non-fine-tuned LLM-based methods by $47\%$ relative improvement but also achieves classification performance comparable to fine-tuned LLMs.
+Our approach enhances GPU memory efficiency by a factor of 188×188$\times$188 × and accelerates both training and inference processes by 26.11×26.$1$.
+Our contributions can be summarized as follows:
+Figure 1.
+The overall architecture of our method for KGC involves categorizing triples based on the type of task, followed by their verbalization.
+This is combined with entity descriptions, which are generated by the subgraph entity description generator, to produce the stimulation prompt.
+Finally, the internal representation (hidden states) of the LLM obtained from the stimulation prompt is used to train a classifier, thereby completing the KGC.
+The only component that needs training in the entire architecture is the data-efficient classifier, making the method easy to train and cheap to deploy.
+  • We propose, to the best of our knowledge, the first framework that directly harnesses the intermediate representation space of non-fine-tuned LLMs for KGC tasks.
+To bridge the semantic gap between the LLM and the KG, we design a subgraph-based entity description generator, ensuring that the model’s internal understanding of entities aligns effectively with KG structures.
+  • We provide an in-depth analysis of why fine-tuned LLMs typically achieve superior KGC performance, identifying the core factors responsible for these gains.
+Building on these insights, we present novel optimization strategies for non-fine-tuned LLMs and are the first to enable them to achieve KGC performance comparable to their fine-tuned counterparts.
+  • We perform extensive experiments across standard KGC datasets to validate the effectiveness and efficiency of our proposed method.
+These substantial improvements highlight the scalability and practical applicability of our method, making it a viable solution for large-scale KGC applications where computational resources and time are critical constraints.
+2. Method
+Our method is proposed in face of the limitations of non-fine-tuned LLMs on KGC.
+In Section 2.2, we present how to precisely extract context-aware representation of knowledge from intermediate layers of LLMs.
+In Section 2.3 we tackle complex relationships in KGs, by the generation of model-friendly entity descriptions through subgraph sampling.
+2.1. Formulation and Notations
 A knowledge graph, $G$, comprises a set of entities, $E$, and relations, $R$.
@@
 We denote $D(e)$, $D(r)$ as the textual description of each entity and each relation.
-We focus on triple classification and relation prediction.
-Triple classification involves evaluating the plausibility of a given knowledge triple, to determine whether the triple is true or not and relation prediction predicts the missing relation in a triple.
-2.2 Stimulating the Intermediate Layers of Large Language Models
-Knowledge graphs consist of numerous triples, and the format of these triples is challenging for the LLM to understand, which may mislead the LLM to retrieve irrelevant knowledge and reveal the wrong responses.
-To enhance the performance of LLMs in KGC, we leverage the internal judgment mechanisms of the language model (Zou et al., 2023).
-Our method involves strategically triggering the necessary depth of knowledge embedded within the LLM to elicit the correct responses required for accurate KGC, thereby improving the overall performance of the KGC task.
-Specifically, since different layers of the language model have different roles (Azaria and Mitchell, 2023), we determine the most suitable layer for the task by experimenting with the validation set.
-For the selection of intermediate layers, see Appendix B.5.
-First, we construct prompt templates to stimulate the model to generate the required classification information in the intermediate layers.
-In the triple classification, our goal is to determine whether the triples in the knowledge graph are true.
-Therefore, we use the following prompt template, $PT_{1}$, to create positive and negative sample pairs for training, as shown in Figure 2.
-Notice that the prompt does not describe positive or negative information for triples.
-Then, we use existing triples in the knowledge graph to construct positive samples $S^{+}$, and employ the negative sampling technique (Lv et al., 2022) to obtain negative samples $S^{-}$.
-Finally, we obtain $N$ training data samples.
-For each training text, we input it into the language model to obtain the representation of each text in the intermediate layers.
-We can obtain information about the plausibility of triples from the hidden state corresponding to the token immediately preceding the model prediction, which is the last token of the input text (Zou et al., 2023).
-Specifically, we capture plausibility information for positive samples $s^{+}\in S^{+}$ by:
-$v\left(s^{+}\right)=\{M_{1:i}\left(s^{+}\right)[-1],0<i<L\},$ (1)
-where $M$ represents the language model, $M_{1:i}$ represents the first $i$ layers of the model, $[-1]$ refers to the hidden state corresponding to the last token, $L$ denotes the number of layers in the model.
-The dimension of $M_{1:i}(s^{+})[-1]$ is the embedding dimension of the language model.
-Negative samples follow the same principle.
-At last, we train a classification model using $v(s^{+})$ and $v(s^{-})$ to distinguish between plausible and implausible triples.
-The goal is to learn a boundary that can separate the hidden states of the intermediate layers of the LLM corresponding to positive samples (which are true) from those corresponding to negative samples (which are false).
-We can use classification models such as multilayer perceptron (MLP), SVM, etc., for classification training.
-Taking logistic regression to determine the plausibility of samples as an example:
-$P\left(Y=\mathbf{1}|x\right)=\frac{1}{1+e^{-\left(w^{T}x+b\right)}},$ (2)
-where x represents $v(s)[l]$, $l$ is the layer index, $w$ and $b$ are weights and bias to be learned.
-For relation prediction, we transform the binary classification into multi-class classification.
-By utilizing the prompt in Figure 7, we stimulate the LLM to generate hidden states that capture the characteristics of different relation types in the KG.
-Details can be found in the Appendix B.1.
-By focusing on extracting classification results directly from the model’s internal processing, we ensure that the responses are both relevant and precise, thereby improving the effectiveness of the KGC task.
-Figure 2: Prompt templates to stimulate the LLM for triple classification.
-2.3 Subgraph Entity Description Generator
-The semantics of entities in the knowledge graph are complex, making it challenging for the language model to interpret based solely on the entity’s name due to inherent ambiguity and polysemy.
-Therefore, we inform the language model about the meaning of each entity in the input.
-The meaning of each entity in KG is generated by its one-hop subgraph.
-Specifically, we extend the prompt template constructed in Section 2.2, aiming to assist the language model in aligning internal entities with entities in the knowledge graph.
-For triple classification, since we employ the probing method to extract KGC features from the samples, we need to ensure that the language model can be stimulated to produce effective hidden states that can be used by the classifier to predict the plausibility of triples.
-Therefore, we construct a new prompt $PT_{2}$, as shown in Figure 2.
-By introducing the $D(e)$, LLMs are able to generate more precise and distinguishable hidden states, thereby improving the overall classification performance.
-To address incomplete entity descriptions within certain datasets, we have devised a method to generate $D(e)$.
-We construct $D(e)$ in two ways.
-The first approach involves a straightforward method where we concatenate all the triples within the one-hop subgraph surrounding the entity $e$, thereby forming $D(e)$ (Zhang et al., 2023).
-However, this direct concatenation can sometimes lead to ineffective or misleading stimuli for the LLM, as the structure of these subgraph triples may significantly differ from the text seen during the LLM’s pre-training phase.
-This divergence can adversely affect the model’s response accuracy.
-To mitigate this issue, our second approach leverages a language model to transform these triples into a format that is more comprehensible to the LLM.
-By rephrasing the triples from the one-hop subgraph around the entity $e$ into natural language text, we produce a more refined and model-friendly version of $D(e)$:
-$D(e)=\operatorname{concat}\{\operatorname{transform}(\{h_{1}\}\{r_{1}\}\{t_{1}\}),$ (3)
-$(h_{1},r_{1},t_{1})\in\operatorname{subgraph}(e)\},$
-where $\operatorname{subgraph(\cdot)}$ is a function that retrieves the set of all triples in the one-hop subgraph where the entity resides; $\operatorname{transform(\cdot)}$ is a function that converts the triple into sentence; $\operatorname{concat(\cdot)}$ is a function to process triple statements.
-For the first approach, $\operatorname{concat(\cdot)}$ is a string concatenation function that joins each transformed triple statement with a separator.
-For the second approach, we need to utilize LLM with the in-context learning technique (Brown et al., 2020) to generate descriptions of entities based on the triples in their subgraph.
-Please refer to Appendix A.1 for details.
-By generating long entity descriptions using the subgraph description generator, we align the internal understanding of entities in the LLM with the entity meanings in the KG, which is similar to the entity linking process in knowledge graphs.
-In this way, the LLM will better understand the semantics of the entities, and improve the KGC performance.
-3 Experiments
+For triple classification, the task is to assess the plausibility of a given knowledge triple, determining whether the triple is true or not.
+As for relation prediction and entity prediction, they aim to predict the missing relation and entity in a triple, respectively.
+2.2. Extracting Knowledge Representations
+Knowledge graphs are composed of numerous triples that encapsulate complex relational and semantic information.
+For non-fine-tuned LLMs, interpreting these intricate structures poses significant challenges, often leading to the retrieval of irrelevant information and the generation of incorrect or misleading responses.
+To enhance the performance of LLMs in KGC, our method leverages the internal judgment mechanisms of LLMs, as shown in (Zou et al., 2023).
+Specifically, we strategically harness the intermediate layers of LLMs to extract rich, context-aware knowledge representations that are pivotal for accurate KGC.
+First, we construct specialized prompt templates designed to stimulate the LLM to generate the necessary classification information within its intermediate layers.
+For triple classification, where the objective is to determine the plautibility of triples in the KG, we employ a prompt template, $PT_{1}$, that creates positive and negative sample pairs for training, as illustrated in Figure 2.
+Notably, the prompt avoids explicitly describing the positivity or negativity of triples.
+Thereafter, we use existing triples in the knowledge graph to construct positive samples $S^{+}$, and employ the negative sampling technique (Lv et al., 2022) to obtain negative samples $S^{-}$.
+This results in a dataset obtaining $N$ training samples.
+Each training sample is input into the LLM to obtain representations from its intermediate layers.
+Specifically, we capture plausibility information and extract hidden states for positive samples $s^{+}\in S^{+}$ as follows:
+(1) $Rep_{j}^{i}\left(s^{+}\right)=\{M_{1:i}\left(s^{+}\right)\left[j\right],0\textless i\textless L,0\textless j\textless N\},$
+where $M$ represents the language model, $M_{1:i}$ represents the first $i$ layers of the model, $[j]$ refers to the hidden state corresponding to the $j$th token, $L$ and $N$ represent the number of layers in the model and the number of tokens resulting from tokenizing $s^{+}$.
+Consistent with findings in (Zou et al., 2023), we utilize the representation of the last token in each input text, as it is more conducive to information extraction due to the autoregressive nature of causal language models.
+Negative samples are processed similarly to ensure consistency in representation extraction.
+The extracted representations $Rep_{j}^{i}(s^{+})$ and $Rep_{j}^{i}(s^{-})$ serve as input features for training a lightweight classifier to distinguish between plausible and implausible triples by learning a decision boundary that effectively separates the hidden states of positive samples from those of negative samples.
+Although the classifiers employed—such as multilayer perceptrons (MLPs) and logistic regression models—are inherently simple, their effectiveness is significantly amplified by the quality and richness of the intermediate layer representations, thereby enhancing the ability to make accurate predictions without the computational overhead associated with fine-tuning the entire LLM.
+For relation prediction, we extend our binary classification framework to a multi-class setting, where the classifier predicts the specific relationship types within the KG.
+Utilizing the $PT_{3}$ shown in Figure 2, we stimulate the LLM to generate hidden states that capture the characteristics of different relation types.
+This approach ensures that the classifier can effectively learn to differentiate between multiple relationship categories from the model’s internal processing, enhancing the accuracy of KGC.
+Figure 2. Prompt templates to stimulate the LLM.
+2.3. Subgraph Entity Description Generator
+The semantics of entities in a KG can be complex and ambiguous, posing significant challenges for LLMs to interpret accurately based solely on entity names.
+To overcome these challenges and enhance the semantic comprehension of entities by LLMs without resorting to computationally intensive fine-tuning, we introduce the subgraph entity description generator, which enriches entity representations by providing additional contextual descriptions derived from their local subgraph structures.
+Specifically, we extend the prompt template constructed in Section 2.2.
+We focus on aligning the language model’s internal entity understanding with those in the knowledge graph.
+Therefore, we construct new prompts $PT_{2}$ and $PT_{4}$, as shown in Figure 2.
+By introducing the entity description $D(e)$, LLMs are able to generate more precise and distinguishable hidden states, thereby improving the overall classification performance.
+Most KGs contain entity descriptions that can be directly used as $D(e)$.
+However, to address incomplete entity descriptions within certain KG datasets, we propose a method to generate $D(e)$.
+Unlike the approach in (Yang et al., 2024), which generates the entity description using only a single triple, our method constructs a more precise $D(e)$ by incorporating subgraph information in two distinct ways.
+For the first approach, we concatenate all the triples within the one-hop subgraph around the entity $e$, thereby forming $D(e)$ (Zhang et al., 2024).
+However, providing information in the form of explicit rules may be less effective for LLM to retrieve and leverage knowledge (Bian et al., 2024), as the structure of these subgraph triples may significantly differ from the text seen during the pre-training phase.
+Therefore, our second approach leverages a language model to transform these triples into a narrative format that is more comprehensible to the LLM.
+By rephrasing the triples from the one-hop subgraph around the entity $e$ into natural language text, we produce a more refined and model-friendly version of $D(e)$ that is more effective for LLMs to retrieve and leverage knowledge.
+Specifically:
+(2) $D(e)=\operatorname{Rephrase}(\{\operatorname{Verbalize}(\{h_{i}\}\{r_{i}\}\{t_{i}\}),$
+$(h_{i},r_{i},t_{i})\in\operatorname{Subgraph}(e)\}),$
+where $\operatorname{Subgraph(\cdot)}$ is a function that automatically retrieves a specified number of triples in the one-hop subgraph where the entity resides; $\operatorname{Verbalize(\cdot)}$ is a function that transform the triple into its textual string, in this work, we use the linear verbalization (Baek et al., 2023); For the first approach, $\operatorname{Rephrase(\cdot)}$ is a string concatenation function that joins each verbalized triple statement with a separator.
+For the second approach, we employ a LLM with the in-context learning technique, then ask it to generate model-friendly descriptions of entities based on the verbalized triples in their subgraph.
+By leveraging the subgraph entity description generator, we align the LLM’s internal entity understanding with those in the KG, akin to entity linking process.
+This alignment not only mitigates ambiguities but also enhances LLM’s semantic comprehension, ultimately boosting KGC performance.
+Method Pred Samples FB13 WN11 15K237N WN18RR UMLS
+TransE S-F ALL 0.815 0.759 0.697 0.884 0.845
+DistMult S-F ALL 0.862 0.871 0.587 0.851 0.864
+ComplEx S-F ALL 0.857 0.862 0.657 0.841 0.871
+RotatE S-F ALL 0.819 0.847 0.685 0.882 0.877
+KG-BERT L-MLP ALL 0.899 0.935 0.560 0.916 0.773
+KGT5 N-L ALL 0.663 0.728 - - -
+LLAMA-7B N-L 0 0.699 0.661 0.573 0.458 0.658
+LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
+KG-LLAMA-7B N-L ALL 0.892 0.955 0.748 0.921 0.858
+KG-LLAMA-SAMPLED N-L 3k/10k 0.727 0.682 0.614 0.574 0.834
+LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
+LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.934 0.862
+LLAMA-DES2 (Tri) H-MLP 3k/10k 0.842 0.755 0.671 0.779 0.782
+LLAMA-DES2 (GPT) H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
+LLAMA-DES-H H-MLP ALL 0.920 - 0.753 0.943 0.863
+Table 1.
+The main experiment results (accuracy) of triple classification.
+In the ‘Pred’ column, $S\mbox{-}F$ uses score function, $L\mbox{-}MLP$ uses MLP on last layer hidden states, $N\mbox{-}L$ uses natural language, and $H\mbox{-}MLP$ uses MLP on intermediate layer hidden states to predict.
+‘Samples’ indicates training sample counts, with $3k/10k$ indicating $10k$ samples for the UMLS dataset and $3k$ for others.
+LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimulus generation; LLAMA-DES-H combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-DES2 (Tri)/(GPT) uses first/second generated descriptions with $PT_{2}$ for stimulus generation.
+Top three results per dataset are bolded.
+3. Experiments
 In the experiments, our objective is to answer the following research questions (RQs):
-  • RQ1: Can frozen LLMs with the proposed probing methods effectively achieve performance on par with their fine-tuned counterparts?
+  • RQ1: Can LLMs with our method effectively achieve performance on par with their fine-tuned counterparts?
   • RQ2: How does our method perform across different language models and classification models?
   • RQ3: What are the data efficiency and computation efficiency of our proposed method for KGC?
-3.1 Datasets
-Dataset |E| |R| # Train # Valid # Test
+3.1. Datasets
+Dataset $|\mathbf{E}|$ $|\mathbf{R}|$ Train # Valid # Test
 FB13 75,043 13 316,232 5,908 23,733
@@
 YAGO3-10 123,182 37 1,079,040 5,000 5,000
-Table 1: Statistical information of datasets.
-In our experiments, we employ five widely-used KG datasets for triple classification: FB13 (Socher et al., 2013), WN11 (Socher et al., 2013), FB15K-237N (Lv et al., 2022), WN18RR (Dettmers et al., 2018), and UMLS (Yao et al., 2019), as well as a commonly used KG dataset for relation prediction: YAGO3-10 (Dettmers et al., 2018).
-For FB13, we use Wikipedia entity descriptions as input sentences.
+Table 2. Statistical information of datasets.
+In our experiments, we employ five widely-used KG datasets for triple classification: FB13 (Socher et al., 2013), WN11 (Socher et al., 2013), FB15K-237N (Lv et al., 2022), WN18RR (Dettmers et al., 2018), and UMLS (Yao et al., 2019).
+For relation prediction, we employ the YAGO3-10 dataset (Dettmers et al., 2018), and for entity prediction, we use WN18RR for evaluation.
+For FB13, we utilize Wikipedia to construct entity descriptions.
 For FB15K-237N, the entity descriptions provided by (Xie et al., 2016) are used.
@@
 In the case of UMLS, entity descriptions provided by the dataset are used.
-In the following, we refer to the entity descriptions as non-generated entity descriptions.
-3.2 Experimental Settings
-To generate entity descriptions, for the first approach in Section 2.3, we concatenate all subgraph triples to generate descriptions.
+In the following, we refer to these entity descriptions as non-generated entity descriptions.
+3.2. Experimental Settings
+To generate entity descriptions, for the first approach in Section 2.3, we concatenate ten randomly selected triples from the one-hop subgraph surrounding each entity to construct the descriptions.
 For the second approach, we employ GPT-3.5-turbo as the LLM.
-Due to cost considerations, we randomly select 4000 test examples from each dataset to generate entity descriptions.
-We mark the results of these methods with an asterisk (*).
 In the following text, we refer to these entity descriptions as generated entity descriptions (Tri)/(GPT).
-We compare our method with four structure information-based methods: TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), ComplEx (Trouillon et al., 2016) and RotatE (Sun et al., 2019), as well as seven additional information-based methods: KG-LLAMA (Yao et al., 2023), KG-BERT (Yao et al., 2019), KG-T5 (Saxena et al., 2022), LLaMA-7b, Alpaca-7B, LLaMA-7B-ICL and Alpaca-7B-ICL (Zhang et al., 2023).
+We compare our method with four structure information-based methods: TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), ComplEx (Trouillon et al., 2016) and RotatE (Sun et al., 2019), as well as seven additional information-based methods: KG-LLAMA (Yao et al., 2024), KG-BERT (Yao et al., 2019), KG-T5 (Saxena et al., 2022), LLaMA-7b, Alpaca-7B, LLaMA-7B-ICL and Alpaca-7B-ICL (Zhang et al., 2024).
 KG-LLAMA, KG-BERT and KGT5 fine-tune the language models using all the training samples.
-LLaMA-7b and Alpaca-7B using the frozen language model to predict.
-LLaMA-7B-ICL and Alpaca-7B-ICL using 2-shot samples with in-context learning to predict.
-Please refer to Section 4.1 for detailed introductions of these methods.
-For structure information-based methods, we reproduce the results with OpenKE (Han et al., 2018).
-For additional information-based methods, we reproduce the results with official code implementation.
-For KGT5, there’s no public record or code for the method on the FB15K237N, WN18RR, UMLS dataset.
-Please refer to Appendix A.2 for detailed experimental settings.
-Method Pred Samples FB13 WN11 15K237N WN18RR UMLS
-TransE S-F ALL 0.815 0.759 0.697 0.884 0.845
-DistMult S-F ALL 0.862 0.871 0.587 0.851 0.864
-ComplEx S-F ALL 0.857 0.862 0.657 0.841 0.908
-RotatE S-F ALL 0.819 0.847 0.685 0.882 0.921
-KG-BERT L-MLP ALL 0.904 0.935 0.560 0.927 0.773
-KGT5 N-L ALL 0.663 0.728 - - -
-LLAMA-7B N-L 0 0.699 0.661 0.573 0.458 0.658
-LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
-KG-LLAMA-7B N-L ALL 0.892 0.955 0.748 0.921 0.858
-KG-LLAMA-SAMPLED N-L 3k/10k 0.727 0.682 0.614 0.574 0.834
-LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
-LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.931 0.862
-LLAMA-DES2 (Tri) H-MLP 3k/10k 0.842 0.755 0.671 0.779 0.782
-LLAMA-DES2 (GPT)* H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
-Table 2: The main experiment results (accuracy) of triple classification.
-In the $Pred$ column, $S\mbox{-}F$ uses score function, $L\mbox{-}MLP$ uses MLP on last layer hidden states, $N\mbox{-}L$ uses natural language, and $H\mbox{-}MLP$ uses MLP on intermediate layer hidden states to predict.
-$Samples$ indicates training sample counts, with $3k/10k$ indicating $10k$ samples for the UMLS dataset and $3k$ for others.
-LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimuli generation; LLAMA-DES-H combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-DES2 (Tri)/(GPT) uses first/second generated descriptions with $PT_{2}$ for stimuli generation.
-Top three results per dataset are bolded.
-3.3 Main Results (For RQ1)
+LLaMA-7B and Alpaca-7B rely on non-fine-tuned LLMs to directly perform predictions.
+LLaMA-7B-ICL and Alpaca-7B-ICL employ LLMs with in-context learning to make predictions (Zhang et al., 2024).
+We give priority to using the results reported in the official baseline paper.
+Please refer to Section 4.1 for detailed introductions of these baseline methods.
+For the negative sampling employed on the datasets, we utilize the same negative sampling technique as (Yao et al., 2019).
+For the intermediate layer selection, we identify the most suitable layer for the task by experimenting with the validation set.
+Each experiment is executed three times with random initializations, and the mean results are reported.
+3.3. Main Results (For RQ1)
 From Table 1, it can be observed that LLAMA-DES-H and LLAMA-DES2 (GPT), when using a partial training set, exhibit comparable performance to KG-LLAMA-7B, the latter being fine-tuned with all training triples.
 Moreover, when trained on subsets of equivalent size, LLAMA-DES-H and LLAMA-DES2 (GPT) markedly surpass KG-LLAMA-SAMPLED, the existing state-of-the-art approach that relies exclusively on LLaMA for KGC.
-For more details, including the results of baseline models such as Alpaca-7B-ICL and Structural-aware-IT, and the performance of our method on the full training set, refer to Appendix B.4.
 The experimental results presented in Table 1 also validate the effectiveness of the approaches proposed in Section 2.2 and Section 2.3.
@@
 Thus, the experimental results substantiate the effectiveness of our proposed methodology.
-Additionally, the frozen LLM can also achieve similar KGC effects as that of fine-tuning the LLM.
+Additionally, the non-fine-tuned LLM can also achieve similar KGC effects as that of fine-tuning the LLM.
 Notice that the performance on the UMLS dataset is special, where the use of generated descriptions hinders the model’s classification accuracy, and the use of non-generated entity descriptions is also worse than not using any entity descriptions.
@@
 Meanwhile, using entity descriptions may be affected by noisy description text, and descriptions generated by GPT-3.5-turbo may be influenced by hallucination, thereby affecting the classification effectiveness of the induced hidden states.
+Meanwhile, our method shows slightly lower performance on the WN11 and FB15K237N datasets compared to KG-LLAMA, which fine-tuned on the full training set.
+This discrepancy may stem from the fact that the pretrained LLM did not acquire knowledge of certain triples in the KG during its pretraining phase, resulting in incorrect predictions even when probing is conducted.
 In addition to the triple classification, we also validate the effectiveness of the proposed probing method on the relation prediction task as shown in Table 3.
 The experimental results similarly demonstrate that the probing method can achieve the nearly same performance as full training set fine-tuning, even when using only 0.6% of the training data.
-Detailed methodology and analysis are provided in Appendix B.1.
-Figure 3: Differences in KGC performance across different layers of the LLAMA-DES2 (GPT) [left] and Mistral-DES2 (GPT) [right].
-The horizontal axis represents the position of the intermediate layers in LLaMA or Mistral (which have different backbones), and the vertical axis represents the dataset.
+For entity prediction, we obtain the plausibility confidence scores of the triples through the classifier and rank them accordingly.
+Table 4 also displays similar results, showing that after using model-friendly descriptions, the LLAMA-DES2 (GPT) method improves the performance of LLAMA-MLP-7B and surpasses its fine-tuned counterparts.
+Figure 3.
+Differences in KGC performance across different layers of the LLAMA-DES2 (GPT) [left] and Mistral-DES2 (GPT) [right].
+The horizontal axis represents the position of the intermediate layers in LLaMA or Mistral, and the vertical axis represents the dataset.
 Method Pred Samples YAGO3-10
@@
 LLaMA-13B N-L 0 0.0040
-KG-LLaMA-7B N-L 0 0.7028
+KG-LLaMA-7B N-L ALL 0.7028
 LLAMA-MLP-7B H-MLP 6996 0.5968
-LLAMA-DES2 (GPT)* H-MLP 6996 0.6824
-Table 3: The experiment results (Hits@1) of relation prediction.
-LLAMA-MLP-7B refers to using LLaMA-7B as the base language model, using MLP as the hidden state classifier, and generating stimuli using $PT_{3}$; LLAMA-DES2 (GPT) refers to using the second generated description combined with $PT_{4}$ to generate stimuli.
-3.4 Ablation Study (For RQ2)
+LLAMA-DES2 (GPT) H-MLP 6996 0.6824
+LLAMA-DES2 (GPT) H-MLP ALL 0.7015
+Table 3. The experiment results (Hits@1) of relation prediction.
+Method Pred Samples WN18RR
+KG-BERT L-MLP ALL 0.1102
+StAR L-MLP ALL 0.2430
+KGT5 N-L ALL 0.1011
+KG-ChatGLM-6B N-L ALL 0.1613
+LLaMA-7B N-L 0 0.0849
+LLaMA-13B N-L 0 0.0991
+KG-LLaMA-7B N-L ALL 0.2415
+LLAMA-MLP-7B H-MLP 10k 0.2108
+LLAMA-DES2 (GPT) H-MLP 10k 0.2433
+LLAMA-DES2 (GPT) H-MLP ALL 0.2495
+Table 4. The experiment results (Hits@1) of entity prediction.
+3.4. Ablation Study (For RQ2)
 Through the analysis depicted in the left portion of Figure 3, it is observed that with an increase in the index of the intermediate layer, the efficacy of hidden states for triple classification demonstrates an initial incline followed by a subsequent decline.
 Intermediate layers closer to natural language text output generally exhibit better prediction performance compared to the lowest layers.
-This observation corroborates the notion posited by Geva et al.
-(2020) that language models encode knowledge differentially across their Feed-Forward layers, with higher layers integrating information from lower layers.
+This observation corroborates the notion posited by (Geva et al., 2020) that language models encode knowledge differentially across their Feed-Forward layers, with higher layers integrating information from lower layers.
 During pretraining, the lower layers may not have stored certain knowledge, and thus are unable to generate hidden states conducive to accurate triple classification.
@@
 This phenomenon may be attributed to the hallucination issue inherent in LLaMA, wherein higher layers could potentially either omit or alter triple information, leading to a blurred distinction between positive and negative triple hidden states induced by stimuli (Jiang et al., 2024; Zou et al., 2023).
-Detailed explanations can be found in Appendix B.5.
 Akin observations are discernible on the right side of Figure 3 for the Mistral model, which has a different backbone from LLaMA.
+This experimental results are consistent with the conclusions of (Jiang et al., 2024), which found that the MLP layers in the later stages have an inhibitory effect on the extracted knowledge, particularly contributing significantly to erroneous output decoding.
+As a result, the middle layers in their experiments also tend to outperform the later layers.
 Meanwhile, as observed in Table 6, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning with only a small sample size.
@@
 The result shows that MLP outperforms all three datasets compared to the other two classification models, while Logistic Regression also exhibits results close to MLP.
-This underscores the efficacy of our stimulus methodology, whereby hidden states of positive and negative samples within a high-dimensional space are easily distinguishable.
+This underscores the efficacy of our stimulus methodology, whereby the model inherently encodes discriminative features for KGC tasks, making hidden states easily distinguishable in high-dimensional space and allowing straightforward classification boundaries to emerge without the need for complex, nonlinear decision functions.
 Method FB13 WN11 FB15K-237N
@@
 LLAMA-MLP 0.855 0.872 0.674
-Table 4: The variation in KGC prediction performance across different datasets using SVM, Logistic Regression, MLP as classification models.
+Table 5.
+The variation in KGC prediction performance across different datasets using SVM, Logistic Regression, MLP as classification models.
 All these methods use LLaMA-7B as the base model, and generate stimuli using $PT_{1}$.
@@
 LLAMA-DES-H 0.901 -
-LLAMA-DES2 (GPT)* 0.912 0.917
+LLAMA-DES2 (GPT) 0.912 0.917
 Mistral KG-MISTRAL-7B 0.891 0.962
@@
 MISTRAL-DES-H 0.880 -
-MISTRAL-DES2 (GPT)* 0.893 0.925
+MISTRAL-DES2 (GPT) 0.893 0.925
 Gemma KG-GEMMA-7B 0.897 0.944
@@
 GEMMA-DES-H 0.891 -
-GEMMA-DES2 (GPT)* 0.867 0.917
-Table 5: The variation in KGC prediction performance across different datasets using LLaMA, Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024) as base models (which have different backbones).
+GEMMA-DES2 (GPT) 0.867 0.917
+Table 6.
+The variation in KGC prediction performance across different datasets using LLaMA, Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024) as base models (which have different backbones).
 Please refer to Table 1 for the number of training samples used by different methods.
-3.5 Efficiency Study (For RQ3)
-As illustrated in Figure 4, our method notably surpasses both KG-LLAMA and KG-BERT on the FB13 when trained with only 100 samples, as opposed to the 3000 samples used by the latter methods.
-For the UMLS, the performance of KG-LLAMA and KG-BERT with 3000 training samples is comparable to that of LLAMA-DES-H trained with only 500 samples.
-Additionally, combining the results from Table 2 and Table 10, for the FB13 dataset, our method attains 97.2% of the performance achievable with the full training set, utilizing only 500 samples, which constitutes a mere 0.079% of the total training samples.
-In contrast, KG-LLAMA and KG-BERT achieve only 78.7% and 63.1% of the maximum performance, respectively.
-For the UMLS, with 9.58% of the training set, our method reaches 90.6% of the full training set performance.
-For other datasets, please refer to Appendix B.6.
-Our experimental results confirm that the language model is capable of generating effective hidden states through the techniques described in Sections 2.2 and 2.3, demonstrating its potential for data-efficient scenarios within the context of KGC tasks.
-Figure 4: The prediction performance of KG-LLAMA, KG-BERT, and LLAMA-DES-H on triple classification varies with the size of the training dataset.
-The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples.
-Meanwhile, we also evaluate the compute efficiency of our method.
-Table 8 shows that our method can achieve significantly lower consumption compared to parameter-efficient fine-tuning even in the setting of the full dataset.
-Compared to fine-tuned LLMs, we boost the GPU memory efficiency to 188×188$\times$188 × during model training and speed up 13.48×13.48$\times$13.48 × overall.
-Please refer to Appendix B.3 for details.
-4 Related Work
-4.1 Knowledge Graph Completion
-Knowledge graphs are typically characterized by their inherent incompleteness, often containing substantial amounts of implicit or missing valuable knowledge (Hogan et al., 2021).
-The prevailing methodologies in KGC can be categorized into two main classes: structure information-based methods and additional information-based methods (Shen et al., 2022b).
-Structure information-based methods leverage the inherent structural data present within knowledge graphs.
-For example, RESCAL (Nickel et al., 2011), DistMult (Yang et al., 2014), COMPGCN (Vashishth et al., 2019), and ConvE (Dettmers et al., 2018) primarily use scoring functions to evaluate the semantic similarities of entity or relation embeddings within a latent embedding space to infer missing information.
-Additional information-based methods incorporate a variety of supplementary data to enrich the knowledge graphs.
-These methods utilize node attributes, entity-related information, or relational path data to enhance the features of knowledge embeddings critical for KGC tasks.
-The prominent examples are KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which represent entities, relations, and triples as textual sequences, turning KGC into a sequence classification problem.
-KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) employ encoder-decoder architecture language models to complete KGC in a generative manner.
-KG-LLAMA (Yao et al., 2023) utilizes triples composed of entity and relation names as sentences, parameter-efficient fine-tuning the LLaMA model for KGC.
-Li et al.
-(2024) enhances KGC with LLM by generating entity descriptions and designing loss functions for fine-tuning.
-4.2 Large Language Models
-Several studies have explored the mechanisms through which Large Language Models process, interpret, and generate factual content.
-As highlighted by Saier et al.
-(2023), these models function as extensive repositories of knowledge, encapsulating a vast array of information about the world within their parameters.
-This insight into LLMs underscores their potential as powerful tools for a wide range of applications, including knowledge graph completion, by leveraging the rich informational content stored in their intricate network structures.
-Ramrakhiyani et al.
-(2023) have conducted evaluations across multiple Large Language Models to assess their knowledge of world geography using probing techniques.
-Similarly, Onoe et al.
-(2022) developed a methodology for automatically constructing cloze prompts by masking out spans, aiming to assess the language models’ understanding of entities through factual probing.
-In addition, Azaria and Mitchell (2023) and Zou et al.
-(2023) have explored training classifiers to estimate the truthfulness of statements based on the activations of the hidden layers in LLMs.
-5 Conclusion
-In this paper, we introduce a novel data-efficient probing method that completes KGC by leveraging the internal knowledge of LLMs.
-Experiments show that using the probing method with generated entity descriptions, frozen LLMs can achieve performance comparable to that of their fine-tuned counterparts.
-Simultaneously, this method saves a significant amount of GPU memory and time compared to parameter-efficient fine-tuning approaches.
-This work explores the potential of frozen LLMs in KGC tasks with both efficacy and efficiency.
-Compared to fine-tuning LLMs, we hit the core of KGC problem in a better way.
-Limitations
-Our current architectural design is only applicable to open-source LLMs or closed-source models that provide access to token embeddings; it is not suitable for closed-source models that exclusively output text.
-Recognizing these limitations, future work will aim to expand our architecture to accommodate closed-source LLMs that only output text or token probabilities.
-Additionally, future efforts could integrate structure information-based KG models, such as TransE, with our methodology in a hybrid approach.
-Appendix A Experimental Details
-A.1 Generate Entity Descriptions using GPT-3.5-turbo
-In Section 2.3, we utilize LLM with the in-context learning technique to generate descriptions for entities based on the triples in their one-hop subgraph.
-In the experiments, we utilize the GPT-3.5 as the LLM.
-The prompt template used to generate entity descriptions is shown in Figure 5.
-Specifically, the $EN(e)$ is the entity name of entity $e$; the $\operatorname{concat\_subgraph()}$ is a function that first converts triples from the one-hop subgraph where the entity is located into sentences, and then concatenates these sentences using a specified delimiter.
-It is important to note that all the triples used in the prompt are from the training set, preventing potential data leakage from the test set in the generated entity descriptions.
-Table 6 shows the descriptions of entities generated using this method, thus achieving "entity alignment" between the knowledge graph and the LLM.
-Entity Generated Description Non-Generated Description
-Gustav Mahler
-Gustav Mahler was a Jewish composer
-known for his significant contributions
-to the world of music. He lived in the
-United States and was of German nationality.
-Gustav Mahler (German; 7 July 1860 – 18 May 1911)
-was an Austro-Bohemian late-Romantic composer,
-and one of the leading conductors of his generation…
-James Hayter
-James Hayter was a male actor born in Scotland,
-known for his profession in acting.
-He passed away in Spain.
-Henry James Goodenough Hayter
-(23 April 1907 – 27 March 1983), better known
-as James Hayter, was a British actor. He is best
-remembered for his roles as Friar Tuck in the film The Story
-of Robin Hood and His Merrie Men (1952) and…
-Table 6: Comparison of generated entity descriptions with non-generated entity descriptions on the FB13 dataset.
-Figure 5: The prompt template used to generate entity descriptions based on the one-hop subgraph where the entity is located for GPT-3.5-turbo.
-A.2 Implementation Details
-Figure 6: The prompt template used in Alpaca-7B-ICL on the WN18RR dataset.
-In our implementation, we experimented with these baselines using the hyperparameters recommended in their paper.
-For the negative sampling employed on the datasets, we utilize the same negative sampling technique as Yao et al.
-(2019) for FB13 and WN11.
-Specifically, FB15K-237N is designed to extract hard negative triples, providing a more challenging evaluation and reducing false negatives in the validation/test datasets during dataset construction.
-A similar approach is adopted for constructing negative samples for WN18RR and UMLS.
-Meanwhile, for all experiments in our study, the training sets contain an equal number of positive and negative samples.
-For the MLP classifier configuration, we opt for a batch size of 64, a learning rate of 3e-5, and utilize the AdamW optimizer, with training restricted to 30 epochs.
-For LLaMA-7B-ICL and Alpaca-7B-ICL, we randomly select one positive sample and one negative sample from the training set of each dataset, respectively.
-The prompt for KGC using Alpaca-7B-ICL on the WN18RR dataset is illustrated in Figure 6.
-Additionally, the hyper-parameters for models trained with the LoRA technique, specifically KG-LLAMA-7B, are outlined in Table 7.
-It’s worth noting that for Table 2, Table 3, Table 4, Table 5 and Figure 4, we determine the most suitable layer for the task by experimenting with the validation set, and then report the results on the test set.
-For Figure 3, the results shown are for different layers of the language model on the test set.
-For the baseline results in Table 2 and Table 3, we give priority to using the results reported in the official baseline paper.
-For results not included in the paper, we replicate the experiments using the hyperparameters recommended by the official sources and conduct multiple attempts and optimizations on these hyperparameters as well.
-Each experiment is executed three times with random initializations, and the mean results are reported.
-Parameter Value
-optimizer adamw
-learning rate 3e-4
-epoch 3
-weight decay 0
-LoRA modules q_proj, k_proj, v_proj, o_proj
-LoRA r 8
-LoRA alpha 16
-LoRA dropout 0.05
-Table 7: Hyper-parameters of KG-LLAMA-7B.
-Appendix B Additional Results and Analysis
-B.1 Relation Prediction
-Figure 7: Prompt templates to stimulate the LLM for relation prediction.
 Method Procedure GPU Memory Time
@@
 (LLM:Probing+MLP)
-12.82G+0.078G
+12.82G+0.018G
 (LLM Parameters+MLP Parameters)
-5h10min+29min+15s
-Table 8: Efficiency Comparsion between KG-LLAMA and LLAMA-MLP on the WN11 dataset. The "Procedure" column represents several key steps of these methods.
-Given an incomplete triple $(h,?,t)$ as query, relation prediction aims to predict the missing relation (denoted as ????).
-For relation prediction, we transform the binary framework of triple classification shown in Figure 1 into a multi-class framework.
-The categories for classification are the relationship types within each knowledge graph.
-Table 3 shows our experimental results on the YAGO3-10 (Dettmers et al., 2018) dataset.
-We use the following prompt template, $PT_{3}$ and $PT_{4}$, to create training samples for classification, as shown in Figure 7.
-$PT_{4}$ introduces the $D(e)$ to generate more precise and distinguishable hidden states, linking entities in the Yago with entities in the LLM, thereby improving the overall relation prediction performance.
-It is worth noting that "ALL" represents 1,079,040 training samples.
-LLaMA-MLP and LLAMA-DES2 (GPT) only use 0.6% of the training set to train the probing classifier.
-It can be seen that LLAMA-DES2 (GPT) with frozen language model parameters achieves results very close to KG-LLAMA-7B, which is fine-tuned on the full dataset.
-B.2 Case Study
-Table 9 shows the prediction results of different models for two samples in the WN11 test set.
-For the first case, it can be seen that among the methods using frozen LLMs, only LLAMA-7B-ICL, LLAMA-MLP, and LLAMA-DES-H predict correctly.
-However, the LLAMA-7B-ICL method predicts "True" for almost all test samples.
-For the second case, it can be seen that only LLAMA-MLP, LLAMA-DES-H, and KG-LLAMA-7B predict correctly.
-However, KG-LLAMA-7B is fine-tuned on the full training set.
-The correct predictions by LLAMA-MLP and LLAMA-DES-H indicate that the LLM itself has the relevant knowledge to make accurate predictions, whereas methods like LLAMA-7B-ICL and ALPACA-7B do not fully utilize the knowledge stored in the language model or generate hallucinations during the response process.
-QUERY MODEL PRED LABEL
-HEAD: parsnip; REL: type of; TAIL: herb; LLAMA-7B N/A TRUE
-LLAMA-7B-ICL TRUE
-ALPACA-7B FALSE
-ALPACA-7B-ICL FALSE
-LLAMA-MLP TRUE
-LLAMA-DES-H TRUE
-KG-LLAMA-7B TRUE
-HEAD: market synset; REL: domain topic; TAIL: brain stem; LLAMA-7B N/A FALSE
-LLAMA-7B-ICL TRUE
-ALPACA-7B TRUE
-ALPACA-7B-ICL TRUE
-LLAMA-MLP FALSE
-LLAMA-DES-H FALSE
-KG-LLAMA-7B FALSE
-Table 9: Case study (triple classification) on two representative examples selected from the WN11 dataset.
-"N/A" indicates that the model’s output does not contain the predicted label information (i.e., yes or no).
-B.3 Efficiency Analysis
-We compared the performance of the KG-LLAMA method and LLAMA-MLP on the WN11 full dataset in terms of GPU memory and time consuming metrics using an Nvidia 3090 GPU.
-The results are shown in Table 8.
-For the computation of GPU memory usage, we use "torch.cuda.memory_allocated" to obtain the peak memory usage at each step.
-For the calculation of GPU memory reduction factor, we focus solely on the training phase, dividing the memory usage of KG-LLAMA by that of LLAMA-MLP.
-For the reduction factor of time consumption, we first sum the time consumed during both the training and inference phases for each method and then divide the results.
-For all methods, the batch size for all procedures is set to 1.
-The KG-LLAMA method employs LoRA for supervised fine-tuning, and its hyperparameters are listed in Table 7, with the epoch set to 1 in the efficiency experiment.
-For the training stage of KG-LLAMA, we employ 8-bit quantization technique, while for the inference stage, we configure not to use beam search.
-For the inference stage of LLAMA-MLP, we probe both the training data and the testing data.
-Thus, the sum of the three times in the table corresponds to the time required for probing the training data, probing the testing data, and the MLP forward pass, respectively.
-It is worth noting that in practical use, our probing method only requires a small portion of the training set to complete the training of classifiers (such as MLP, Logistic Regression, etc.).
-Here, we used the full dataset for fairness in testing.
-We can clearly observe in the Table 8 that the proposed probing method, which uses frozen language models, consumes less GPU memory at all stages compared to KG-LLAMA, even though KG-LLAMA employs parameter-efficient fine-tuning techniques like LoRA to save memory.
-Additionally, since our method only requires a forward pass through the LLM, it consumes significantly less time compared to KG-LLAMA.
-Method Pred Samples FB13 WN11 15K237N WN18RR UMLS
-LLAMA-7B N-L 0 0.699 0.661 0.573 0.458 0.658
-ALPACA-7B N-L 0 0.226 0.696 0.561 0.714 0.526
-LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
-ALPACA-7B-ICL N-L 2 0.675 0.706 0.607 0.691 0.605
-STRUCTURAL-AWARE-IT N-L ALL 0.907 0.959 0.752 0.927 0.881
-ALPACA-DES N-L 0 0.611 - 0.590 0.548 0.525
-LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
-LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.931 0.862
-LLAMA-DES2 (GPT)* H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
-LLAMA-DES-H H-MLP ALL 0.905 - 0.742 0.937 0.866
-Table 10: The experiment results (accuracy) of triple classification. ALPACA-DES combines non-generated descriptions with $PT_{2}$ for Alpaca-7B.
-B.4 Additional Baselines
-From Table 10, it is noteworthy that ALPACA-DES fails to outperform ALPACA-7B, suggesting that mere incorporation of entity descriptions may not yield direct performance enhancements, considering the possibility of LLMs revealing the wrong responses and being influenced by hallucinations.
-Meanwhile, in our study, we did not compare our method with KOPA because it integrates KG models (such as TransE/RotatE) with language models (like LLaMA) in a hybrid approach, thereby blending multiple approaches rather than solely utilizing the capabilities of LLMs (Zhang et al., 2023).
-In contrast, our current research exclusively focuses on the utilization of language models (like LLaMA), rendering KOPA an unsuitable baseline for our purposes.
-Nonetheless, we reproduced the Structural-aware-IT model proposed by Zhang et al.
-(2023), which incorporates in-context learning and instruction tuning.
-This model samples the neighborhoods of the head entity $h$ and tail entity $t$, including the textual descriptions of neighboring triples in the demonstration prompt, followed by fine-tuning on the entire training set.
-Like KOPA, this approach integrates structural information from the KG with LLM capabilities.
-Despite this, our method, which exclusively utilizes LLMs without integrating KG structural information, achieves performance comparable to the Structural-aware-IT model, which combines both KG structure and LLM capabilities.
-B.5 Analysis of the Performance in the Intermediate Layers of LLM
-The selection of the optimal layer index in LLMs serves as a hyperparameter, which we may select based on empirical evidence or adjust as needed when a validation set is available.
-Meanwhile, our experiments found a strong consistency for the same LLM across different knowledge graphs (for example, for LLAMA, the 15th layer often yields the best KGC performance), as shown in Figure 3, which reduces the complexity of utilization.
-Additionally, the superior performance of the intermediate layers of language models, as compared to those closer to the output layer, may be attributed to the potential for each layer of the model to experience hallucinations.
-Moreover, each layer of a language model can potentially propagate hallucinations from previous layers, manifesting as deliberate falsehoods or the expression of misleading information (Zou et al., 2023).
-This phenomenon can lead to scenarios where models often know the correct answer even while generating incorrect outputs (Zou et al., 2023).
-Concurrently, our experimental results are consistent with the conclusions of Jiang et al.
-(2024), which found that the MLP layers in the later stages have an inhibitory effect on the extracted knowledge, particularly contributing significantly to erroneous output decoding.
-As a result, the middle layers in their experiments also tend to outperform the later layers.
-B.6 Additional Analysis of the Data Efficiency
-From Figure 8, we can observe the following: For the FB13 dataset, using 400 training samples, which is 0.06% of the training set, achieves 97.0% of the performance of using the entire training set; For the FB15K-237N dataset, using 0.57% of the training set achieves 99.6% of the performance of using the entire training set; For the WN18RR dataset, using 0.46% of the training set achieves 98.8% of the performance of using the entire training set; For the UMLS dataset, using 9.58% of the dataset achieves 90.6% of the performance of using the entire training set.
-Figure 8: The LLAMA-DES-H prediction performance on KGC varies with the size of the training dataset.
-The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples, and the vertical axis represents the accuracy on the test set.
-It can be observed that the LLAMA-DES-H method exhibits slightly lower data efficiency on the UMLS dataset compared to other methods.
-This phenomenon may be attributed to the fact that UMLS is a domain-specific knowledge graph, whereas the other datasets are encyclopedic knowledge graphs.
-The increased complexity of the decision boundary in the hidden states of the positive and negative example triplets generated by LLaMA necessitates a larger volume of data to adequately model the boundary.
-B.7 Analysis of Methods for Entity Description Generation
-From Table 2, it can be observed that the performance of LLAMA-DES2 (Tri) is significantly inferior to that of LLAMA-DES2 (GPT), and even LLAMA-MLP.
-This indicates that using subgraph triples as entity descriptions directly does not enhance the model’s classification performance; it may even impair its judgment.
-This phenomenon may be attributed to subgraph triples diverging significantly from the language model’s pre-training corpus, thus providing erroneous stimulus information.
-However, organizing triples into coherent semantic natural language text using methods in Appendix A.1 appears to mitigate this impact.
-Figure 9: We visualize the hidden states obtained from stimulation in the FB13 test set in three-dimensional space using PCA.
-The high dimensional vectors are obtained using the LLAMA-DES2 (GPT) method, collecting the language model’s hidden states from the 16th layer through stimulation.
-B.8 Analysis of Using Frozen LLM in KGC
-It has been observed that the predictive performance achieved by the frozen language models can not achieve comparable performance with supervised fine-tuning the language models (Zhang et al., 2023; Yao et al., 2023).
-But the knowledge graphs used in the experiments are encyclopedic knowledge graphs, such as FB15K-237 (Toutanova and Chen, 2015), which are subsets of FB15k (Bordes et al., 2013) derived from Wikipedia.
-LLMs like LLaMA (Touvron et al., 2023), which are pretrained on datasets such as Wikipedia, CommonCrawl, etc., have stored this type of world knowledge (Davison et al., 2019).
-Liu et al.
-(2023) shows that prompting can effectively access the knowledge within LLMs without fine-tuning.
-Therefore, frozen LLMs have sufficient knowledge and capability to complete KGC tasks.
+2h44min+15s
+(Probing+MLP Forward)
+Table 7. Efficiency Comparsion between KG-LLAMA-7B and LLAMA-MLP on the WN11 full dataset. The ‘Procedure’ column represents several key steps of these methods.
+3.5. Efficiency Study (For RQ3)
+As illustrated in Figure 4, our method outperforms both KG-LLAMA and KG-BERT on the FB13 when trained with only 100 samples, as opposed to the 3000 samples used by the latter methods.
+For the UMLS, the performance of KG-LLAMA and KG-BERT with 3000 training samples is comparable to that of LLAMA-DES-H trained with only 500 samples.
+Additionally, combining the results from Table 1, for the FB13 dataset, our method attains 97.2% of the performance achievable with the full training set, utilizing only 500 samples, which constitutes a mere 0.079% of the total training samples.
+In contrast, KG-LLAMA and KG-BERT achieve only 78.7% and 63.1% of the maximum performance, respectively.
+For the UMLS, with 9.58% of the training set, our method reaches 90.6% of the full training set performance.
+Our experimental results valid that the language model is capable of generating effective hidden states through our method, demonstrating its potential for data-efficient scenarios within the context of KGC tasks.
+In addition to data efficiency, we also evaluate the computational efficiency of our method.
+Table 7 shows that our method achieves significantly reduced resource usage compared to parameter-efficient fine-tuning.
+For the computation of GPU memory usage, we use ”torch.cuda.memory_allocated” to obtain the peak memory usage at each step.
+The KG-LLAMA method employs LoRA and 8-bit quantization technique for supervised fine-tuning.
+Compared to fine-tuned 7B LLMs, we boost the GPU memory efficiency to 188×188$\times$188 × during model training and speed up 26.11×26.$1$ overall.
+Notably, the training overhead of our method lies solely in the classifier, which is independent of the LLM size.
+Therefore, when scaling LLMs to larger sizes, such as 70B, the substantial overhead associated with backpropagation in fine-tuning becomes increasingly burdensome.
+In these scenarios, the advantages of our method in terms of GPU memory usage and time consumption become even more pronounced.
+Figure 4.
+The prediction performance of KG-LLAMA, KG-BERT, and LLAMA-DES-H on triple classification varies with the size of the training dataset.
+The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples.
+Figure 5. Generalization across Datasets, LLAMA-DES-H. The horizontal axis represents the training set, and the vertical axis represents the test set.
+Figure 6. Performance of LLAMA-DES2 (GPT) versus the number of sampled triplets.
+4. Related Work
+4.1. Knowledge Graph Completion
+Knowledge graphs are typically characterized by their inherent incompleteness (Hogan et al., 2021).
+The prevailing methodologies in KGC can be categorized into two main classes: structure information-based methods and additional information-based methods (Shen et al., 2022b).
+Structure information-based methods leverage the inherent structural data present within knowledge graphs.
+For example, RESCAL (Nickel et al., 2011), DistMult (Yang et al., 2014), COMPGCN (Vashishth et al., 2019), and ConvE (Dettmers et al., 2018) primarily use scoring functions to evaluate the semantic similarities of entity or relation embeddings within a latent embedding space to infer missing information.
+Additional information-based methods incorporate a variety of supplementary data to enrich the knowledge graphs.
+The prominent examples are KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which represent entities, relations, and triples as textual sequences, turning KGC into a sequence classification problem.
+KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) employ encoder-decoder architecture language models to complete KGC in a generative manner.
+And (Li et al., 2024) enhances KGC with LLM by generating entity descriptions and designing loss functions for fine-tuning.
+More recently, KG-LLAMA (Yao et al., 2024) is the first to adopt a decoder-only LLM architecture for KGC by formulating the task as question answering through an instruction-tuning-based prompt, followed by parameter-efficient fine-tuning on LLaMA.
+Notably, KG-LLAMA leverages only LLaMA’s own capabilities without incorporating additional hybrid approaches, which aligns with the core principle of our proposed framework.
+4.2. Large Language Models
+Several studies have explored the mechanisms through which Large Language Models process, interpret, and generate factual content.
+As highlighted by (Saier et al., 2023), these models function as extensive repositories of knowledge, encapsulating a vast array of information about the world within their parameters.
+This insight into LLMs underscores their potential as powerful tools for a wide range of applications, including knowledge graph completion, by leveraging the rich informational content stored in their intricate network structures.
+(Ramrakhiyani et al., 2023) have conducted evaluations across multiple Large Language Models to assess their knowledge of world geography using probing techniques.
+Similarly, (Jin et al., 2024) used probing to measure the abilities of different LLM layers to learn various levels of concepts.
+In addition, (Liu et al., 2024) investigated whether a universal truthfulness hyperplane exists within LLMs by employing probing methods.
+5. Conclusion
+This paper proposes a novel framework that leverages the latent representation space of LLMs to improve KGC.
+By integrating subgraph-based entity descriptions, our approach bridges the semantic gap between LLMs and KGs, enabling LLMs to achieve performance on par with their fine-tuned counterparts.
+Simultaneously, our method enhances data efficiency and saves a significant amount of GPU memory and time compared to parameter-efficient fine-tuning.
+This work explores the potential of LLMs in KGC tasks with both efficacy and efficiency.
+Compared to fine-tuning, we hit the core of KGC problem in a better way.

### v4

@@
-Bridging LLMs and KGs without Fine-Tuning: Intermediate Probing Meets Subgraph-Aware Entity Descriptions
+Unlock the Power of Frozen LLMs in Knwoledge Graph Completion
 Abstract.
@@
 However, direct fine-tuning of LLMs for KGC, though effective, imposes substantial computational and memory overheads, while utilizing non-fine-tuned LLMs is efficient but yields suboptimal performance.
-In this work, we propose a novel framework that synergizes the strength of LLMs with robust knowledge representation to enable effective and efficient KGC.
+In this work, we propose a novel framework that synergizes the strengths of LLMs with robust knowledge representation to enable effective and efficient KGC.
 We extract the context-aware hidden states of knowledge triples from the intermediate layers of LLMs, thereby capturing rich semantic and relational nuances.
-These represenation are then utilized to train a data-effecient classifier tailored specifically for KGC tasks.
-To bridge the semanitic gaps between LLMs and KGs, we employ subgraph sampling on KGs to generate model-friendly entity descriptions, ensuring that the representations are comprehensive and well aligned with the KG structure.
+These representations are then utilized to train a data-efficient classifier tailored specifically for KGC tasks.
+To bridge the semantic gaps between LLMs and KGs, we employ subgraph sampling on KGs to generate model-friendly entity descriptions.
+We further adopt sliced mutual information (SMI) as a principled metric to quantify the task-specific information encoded in these representations.
 Extensive experiments on standard benchmarks validate the efficiency and effectiveness of our approach.
-We achieve a 47% relative improvement over previous methods based on non-fine-tuned LLMs and, to our knowledge, are the first to achieve classification performance comparable to fine-tuned LLMs while enhancing GPU memory efficiency by 188×188$\times$188 × and accelerating training and inference by 26.11×26.$1$.
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
+We achieve a 47% relative improvement over previous methods based on non-fine-tuned LLMs and, to our knowledge, are the first to achieve classification performance comparable to fine-tuned LLMs while enhancing GPU memory efficiency by $188\times$ and accelerating training and inference by $26.11\times$.
+Knowledge Graph Completion, Large Language Models
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: CIKM; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Reasoning about belief and knowledgeccs: Computing methodologies Information extractionccs: Computing methodologies Neural networksccs: Computing methodologies Supervised learning by classification
 1. Introduction
-Knowledge graph completion has become a crucial endeavor to improve the application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b).
-Traditional knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which is vulnerable to the sparseness of KGs (Yao et al., 2019).
-By leveraging the powerful contextual modeling capabilities and extensive stored world knowledge of LLMs, we can effectively mitigate these challenges, and enhance the performance of KGC (Wei et al., 2024).
-These approaches generally fall into two types: using non-fine-tuned LLMs with prompting techniques and fine-tuned LLMs specifically for KGC.
-The first type leverages LLMs with prompt techniques, such as in-context learning.
-These techniques enable models to perform KGC directly through natural language outputs (Yang et al., 2024; Wei et al., 2024; Yao et al., 2024), harnessing the pre-trained knowledge embedded within these models without additional training.
-While this approach offers computational efficiency, it may not fully capture the task-specific nuances required for optimal KGC performance.
-In contrast, the latter type involves fine-tuning the LLMs using supervised techniques to adapt models specifically for KGC tasks.
-This method involves fine-tuning LLMs, such as LLaMA, through the use of training samples (Yao et al., 2024), which can lead to improved performance by aligning the models more closely with the KGC objectives (Zhang et al., 2024; Yao et al., 2024).
-However, fine-tuning demands significant increases in time and resource expenditure.
-This dichotomy between computational efficiency and task-specific performance raises a fundamental question: How can we harness LLMs for effective and efficient KGC without incurring the high costs associated with fine-tuning?
-Answering this question necessitates a deeper look at the limitations of LLMs compared to their fine-tuned counterparts.
-Non-fine-tuned LLMs, while cost-effective, risk generating misleading outputs in the form of hallucinations (Zhao et al., 2023) and underutilizing their latent knowledge (Wang et al., 2023).
-Moreover, the intricate relational structures and semantic subtleties embedded in KGs compound the challenge, since LLMs are not inherently designed to parse or disambiguate complex entity linkages, without explicitly fine-tuning on such tasks (Li et al., 2024).
-In this paper, we propose a novel framework that reconciles these competing demands by fully exploiting the knowledge representation capabilities of LLMs without fine-tuning, as shown in Figure 1.
-Our method strategically extracts intermediate hidden states from LLMs to capture rich, context-aware representations of knowledge triples, which are then used to train a lightweight, data-efficient classifier specifically tailored for KGC.
-To bridge the gap between LLMs and KGs, we incorporate an innovative subgraph sampling technique to generate model-friendly entity descriptions, enabling the LLMs to better interpret and disambiguate relational patterns to further enhance knowledge representation.
-By focusing on these intermediate LLM representations, our framework mitigates hallucination risks while retaining the benefits of a non-fine-tuned paradigm, allowing for both high accuracy and reduced computational costs.
-Extensive experiments demonstrate that our framework not only outperforms existing non-fine-tuned LLM-based methods by $47\%$ relative improvement but also achieves classification performance comparable to fine-tuned LLMs.
-Our approach enhances GPU memory efficiency by a factor of 188×188$\times$188 × and accelerates both training and inference processes by 26.11×26.$1$.
-Our contributions can be summarized as follows:
 Figure 1.
@@
 The only component that needs training in the entire architecture is the data-efficient classifier, making the method easy to train and cheap to deploy.
-  • We propose, to the best of our knowledge, the first framework that directly harnesses the intermediate representation space of non-fine-tuned LLMs for KGC tasks.
-To bridge the semantic gap between the LLM and the KG, we design a subgraph-based entity description generator, ensuring that the model’s internal understanding of entities aligns effectively with KG structures.
-  • We provide an in-depth analysis of why fine-tuned LLMs typically achieve superior KGC performance, identifying the core factors responsible for these gains.
-Building on these insights, we present novel optimization strategies for non-fine-tuned LLMs and are the first to enable them to achieve KGC performance comparable to their fine-tuned counterparts.
-  • We perform extensive experiments across standard KGC datasets to validate the effectiveness and efficiency of our proposed method.
-These substantial improvements highlight the scalability and practical applicability of our method, making it a viable solution for large-scale KGC applications where computational resources and time are critical constraints.
+Knowledge graph completion has become a crucial endeavor to improve the application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b).
+Traditional knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which is vulnerable to the sparseness of KGs (Yao et al., 2019).
+Leveraging the powerful contextual modeling capabilities and the implicit world knowledge encoded in LLMs, recent works have demonstrated the potential to effectively mitigate these issues and enhance KGC performance (Wei et al., 2024).
+These approaches generally fall into two types: using non-fine-tuned LLMs with prompting techniques and fine-tuned LLMs specifically for KGC.
+The first type leverages LLMs with prompt techniques, such as in-context learning.
+These techniques enable models to perform KGC directly through natural language outputs (Yang et al., 2024; Wei et al., 2024; Yao et al., 2025), harnessing the pre-trained knowledge embedded within these models without additional training.
+While this approach offers computational efficiency, it may not fully capture the task-specific nuances required for optimal KGC performance.
+In contrast, the latter type involves fine-tuning the LLMs using supervised techniques to adapt models specifically for KGC tasks.
+This method involves fine-tuning LLMs, such as LLaMA, through the use of training samples (Yao et al., 2025), which can lead to improved performance by aligning the models more closely with the KGC objectives (Zhang et al., 2024; Yao et al., 2025).
+However, fine-tuning demands significant increases in time and resource expenditure.
+This dichotomy between computational efficiency and task-specific performance raises a fundamental question: How can we harness LLMs for effective and efficient KGC without incurring the high costs associated with fine-tuning?
+Answering this question necessitates a deeper understanding of why fine-tuned LLMs outperform their non-fine-tuned counterparts.
+Non-fine-tuned LLMs, while cost-effective, risk generating misleading outputs in the form of hallucinations (Zhao et al., 2023) and underutilizing their latent knowledge (Wang et al., 2023).
+Moreover, the intricate relational structures and semantic subtleties embedded in KGs compound the challenge, since LLMs are not inherently designed to parse or disambiguate complex entity linkages (Li et al., 2024).
+Crucially, fine-tuning mitigates hallucinations (Zhou et al., 2024) and aligns LLM representations with KG semantics through supervised parameter updates.
+Motivated by this insight, we propose a novel framework that reconciles these competing demands by fully exploiting the knowledge representation capabilities of LLMs without fine-tuning, as shown in Figure 1.
+Our method strategically extracts intermediate hidden states from LLMs to capture rich, context-aware representations of knowledge triples, which are then used to train a lightweight, data-efficient classifier specifically tailored for KGC.
+To better align LLMs with KG semantics, we introduce a novel subgraph sampling technique that generates model-friendly entity descriptions, facilitating the interpretation and disambiguation of relational patterns and thereby enhancing knowledge representation.
+Furthermore, we introduce Sliced Mutual Information (SMI) as a principled information-theoretic measure to quantify the task-specific information encoded within these representations.
+By systematically integrating these techniques, our framework effectively reduces hallucinations and semantic misalignment commonly observed in frozen LLMs, achieving accuracy comparable to fine-tuned counterparts while significantly reducing computational overhead.
+Extensive experiments demonstrate that our framework not only outperforms existing non-fine-tuned LLM-based methods by $47\%$ relative improvement but also achieves classification performance comparable to fine-tuned LLMs.
+Our approach enhances GPU memory efficiency by a factor of $188\times$ and accelerates both training and inference processes by $26.11\times$.
+Our contributions can be summarized as follows:
+  • We propose, to the best of our knowledge, the first framework that directly leverages the intermediate representation space of frozen LLMs for KGC tasks.
+To facilitate semantic alignment between LLM representations and KG entities, we introduce a novel subgraph-based entity description generator, providing LLMs with rich, model-friendly textual contexts to effectively bridge the semantic gap.
+  • We provide an in-depth analysis to reveal the underlying reasons behind the superior performance of fine-tuned LLMs.
+We further employ SMI to quantify task-specific information encoded in LLM representations.
+Guided by these insights, we propose novel optimization strategies that, for the first time, enable frozen LLMs to match the KGC performance of their fine-tuned counterparts.
+  • We conduct comprehensive experiments across standard KGC datasets to validate the effectiveness and efficiency of our proposed method.
+These substantial improvements highlight the scalability and practicality of our method, making it a viable solution for large-scale KGC applications under resource-constrained scenarios.
 2. Method
@@
 In Section 2.3 we tackle complex relationships in KGs, by the generation of model-friendly entity descriptions through subgraph sampling.
+Furthermore, in Section 2.4, we employ SMI as a principled information-theoretic metric to quantify the task-specific information encoded within the extracted representations, providing deeper insights into their effectiveness.
 2.1. Formulation and Notations
@@
 It can be represented as a set of triples $G=\{(h,r,t)\}$, where $h\in E$ is the head entity, $t\in E$ is the tail entity, we collectively refer to the head entity and tail entity as $e$, and $r\in R$ represents the relation between them.
-$D$ is the set of descriptions for each entity and relation.
-We denote $D(e)$, $D(r)$ as the textual description of each entity and each relation.
+$D$ is the set of long descriptions for each entity and relation.
+We denote $D(e)$, $D(r)$ as the long textual description of each entity and each relation (hereafter simply referred to as descriptions).
+Additionally, we define $EN(e)$ as the entity name of an entity $e$.
+For instance, given the entity ‘00952182’, its entity name is $EN(\text{`00952182'})=\text{`voice'}$, and its description $D(\text{`00952182'})$ can be ‘utter with vibrating vocal chords’.
 For triple classification, the task is to assess the plausibility of a given knowledge triple, determining whether the triple is true or not.
@@
 2.2. Extracting Knowledge Representations
-Knowledge graphs are composed of numerous triples that encapsulate complex relational and semantic information.
-For non-fine-tuned LLMs, interpreting these intricate structures poses significant challenges, often leading to the retrieval of irrelevant information and the generation of incorrect or misleading responses.
-To enhance the performance of LLMs in KGC, our method leverages the internal judgment mechanisms of LLMs, as shown in (Zou et al., 2023).
-Specifically, we strategically harness the intermediate layers of LLMs to extract rich, context-aware knowledge representations that are pivotal for accurate KGC.
+Figure 2. Prompt templates to stimulate the LLM.
+Knowledge graphs comprise numerous triples encoding complex relational and semantic information.
+However, directly prompting pre-trained, frozen LLMs, despite their extensive world knowledge, frequently results in hallucinations or incorrect responses.
+Although fine-tuning effectively mitigates this issue by optimizing the model’s parameters to produce KG-consistent outputs, it incurs significant computational overhead.
+Motivated by recent findings (Zou et al., 2023; Orgad et al., 2024) that accurate relational knowledge is often internally encoded within LLMs yet not reliably reflected in their outputs, we propose directly probing intermediate layer representations as a principled alternative to fine-tuning.
+This approach leverages the internal semantic encoding capabilities of frozen LLMs, thereby mitigating the computational overhead and hallucination problems associated with direct textual generation.
 First, we construct specialized prompt templates designed to stimulate the LLM to generate the necessary classification information within its intermediate layers.
-For triple classification, where the objective is to determine the plautibility of triples in the KG, we employ a prompt template, $PT_{1}$, that creates positive and negative sample pairs for training, as illustrated in Figure 2.
+For triple classification, where the objective is to determine the plausibility of KG triples, we employ a prompt template $PT_{1}$ to construct positive and negative sample pairs for training, as illustrated in Figure 2.
 Notably, the prompt avoids explicitly describing the positivity or negativity of triples.
-Thereafter, we use existing triples in the knowledge graph to construct positive samples $S^{+}$, and employ the negative sampling technique (Lv et al., 2022) to obtain negative samples $S^{-}$.
-This results in a dataset obtaining $N$ training samples.
-Each training sample is input into the LLM to obtain representations from its intermediate layers.
+Subsequently, we construct positive samples $S^{+}$ from existing KG triples, and employ the negative sampling technique (Yao et al., 2019) to generate negative samples $S^{-}$.
+This process results in a training dataset containing $N$ samples.
+Each training sample is tokenized and fed into the LLM to obtain representations from its intermediate layers.
 Specifically, we capture plausibility information and extract hidden states for positive samples $s^{+}\in S^{+}$ as follows:
-(1) $Rep_{j}^{i}\left(s^{+}\right)=\{M_{1:i}\left(s^{+}\right)\left[j\right],0\textless i\textless L,0\textless j\textless N\},$
-where $M$ represents the language model, $M_{1:i}$ represents the first $i$ layers of the model, $[j]$ refers to the hidden state corresponding to the $j$th token, $L$ and $N$ represent the number of layers in the model and the number of tokens resulting from tokenizing $s^{+}$.
+(1) $Rep_{j}^{l}\left(s^{+}\right)=\{M_{1:l}\left(s^{+}\right)\left[j\right],0\textless l\textless L,0\textless j\textless N\},$
+where $M$ represents the language model, $M_{1:l}$ represents the first $l$ layers of the model, $[j]$ refers to the hidden state corresponding to the $j$th token, $L$ and $N$ represent the number of layers in the model and the number of tokens resulting from tokenizing $s^{+}$, respectively.
 Consistent with findings in (Zou et al., 2023), we utilize the representation of the last token in each input text, as it is more conducive to information extraction due to the autoregressive nature of causal language models.
 Negative samples are processed similarly to ensure consistency in representation extraction.
-The extracted representations $Rep_{j}^{i}(s^{+})$ and $Rep_{j}^{i}(s^{-})$ serve as input features for training a lightweight classifier to distinguish between plausible and implausible triples by learning a decision boundary that effectively separates the hidden states of positive samples from those of negative samples.
+The extracted representations $Rep_{j}^{l}(s^{+})$ and $Rep_{j}^{l}(s^{-})$ serve as input features to train a lightweight classifier.
+This classifier distinguishes plausible from implausible triples by learning a decision boundary that effectively separates positive and negative samples in the representation space.
 Although the classifiers employed—such as multilayer perceptrons (MLPs) and logistic regression models—are inherently simple, their effectiveness is significantly amplified by the quality and richness of the intermediate layer representations, thereby enhancing the ability to make accurate predictions without the computational overhead associated with fine-tuning the entire LLM.
@@
 This approach ensures that the classifier can effectively learn to differentiate between multiple relationship categories from the model’s internal processing, enhancing the accuracy of KGC.
-Figure 2. Prompt templates to stimulate the LLM.
+Figure 3. Prompt used for generating model-friendly entity descriptions via LLM rephrasing. In-context examples omitted for brevity.
 2.3. Subgraph Entity Description Generator
-The semantics of entities in a KG can be complex and ambiguous, posing significant challenges for LLMs to interpret accurately based solely on entity names.
-To overcome these challenges and enhance the semantic comprehension of entities by LLMs without resorting to computationally intensive fine-tuning, we introduce the subgraph entity description generator, which enriches entity representations by providing additional contextual descriptions derived from their local subgraph structures.
-Specifically, we extend the prompt template constructed in Section 2.2.
-We focus on aligning the language model’s internal entity understanding with those in the knowledge graph.
+Entities in a KG often exhibit semantic ambiguity, posing significant challenges for frozen LLMs that rely solely on entity names.
+Fine-tuning alleviates this ambiguity by aligning LLM entity representations with KG entities.
+However, for frozen LLMs, achieving such semantic alignment without parameter updates remains challenging.
+To bridge this gap, we propose a principled subgraph-besed approach that generates rich, context-aware entity descriptions from local subgraph contexts.
+These descriptions provide rich semantic context, enabling frozen LLMs to achieve semantic alignment comparable to fine-tuned models, without costly parameter updates.
+Specifically, we achieve this by extending the prompt template constructed in Section 2.2.
+We focus on aligning the language model’s internal entity understanding with those in the KG.
 Therefore, we construct new prompts $PT_{2}$ and $PT_{4}$, as shown in Figure 2.
-By introducing the entity description $D(e)$, LLMs are able to generate more precise and distinguishable hidden states, thereby improving the overall classification performance.
-Most KGs contain entity descriptions that can be directly used as $D(e)$.
-However, to address incomplete entity descriptions within certain KG datasets, we propose a method to generate $D(e)$.
-Unlike the approach in (Yang et al., 2024), which generates the entity description using only a single triple, our method constructs a more precise $D(e)$ by incorporating subgraph information in two distinct ways.
-For the first approach, we concatenate all the triples within the one-hop subgraph around the entity $e$, thereby forming $D(e)$ (Zhang et al., 2024).
-However, providing information in the form of explicit rules may be less effective for LLM to retrieve and leverage knowledge (Bian et al., 2024), as the structure of these subgraph triples may significantly differ from the text seen during the pre-training phase.
-Therefore, our second approach leverages a language model to transform these triples into a narrative format that is more comprehensible to the LLM.
-By rephrasing the triples from the one-hop subgraph around the entity $e$ into natural language text, we produce a more refined and model-friendly version of $D(e)$ that is more effective for LLMs to retrieve and leverage knowledge.
-Specifically:
-(2) $D(e)=\operatorname{Rephrase}(\{\operatorname{Verbalize}(\{h_{i}\}\{r_{i}\}\{t_{i}\}),$
-$(h_{i},r_{i},t_{i})\in\operatorname{Subgraph}(e)\}),$
-where $\operatorname{Subgraph(\cdot)}$ is a function that automatically retrieves a specified number of triples in the one-hop subgraph where the entity resides; $\operatorname{Verbalize(\cdot)}$ is a function that transform the triple into its textual string, in this work, we use the linear verbalization (Baek et al., 2023); For the first approach, $\operatorname{Rephrase(\cdot)}$ is a string concatenation function that joins each verbalized triple statement with a separator.
-For the second approach, we employ a LLM with the in-context learning technique, then ask it to generate model-friendly descriptions of entities based on the verbalized triples in their subgraph.
-By leveraging the subgraph entity description generator, we align the LLM’s internal entity understanding with those in the KG, akin to entity linking process.
-This alignment not only mitigates ambiguities but also enhances LLM’s semantic comprehension, ultimately boosting KGC performance.
+Most KGs do not contain explicit entity descriptions that can be directly used as $D(e)$, or provide only limited and incomplete descriptions.
+Therefore, we propose a method to generate comprehensive and informative entity descriptions $D(e)$.
+Unlike the approach in (Yang et al., 2024), which generates the entity description using only a single triple, our method constructs a more comprehensive and precise $D(e)$ by incorporating richer subgraph contexts in two distinct ways.
+For the first approach, following (Zhang et al., 2024), we construct $D_{1}(e)$ by concatenating the verbalized triples within the one-hop subgraph around the entity $e$:
+(2) $D_{1}(e)=\operatorname{CONCAT}\big{(}\{\operatorname{Verbalize}(EN(h_{i}),EN(r_{i}),EN(t_{i})),$
+$(h_{i},r_{i},t_{i})\in\operatorname{Subgraph}(e)\}\big{)},$
+where $\operatorname{Subgraph(\cdot)}$ is a function that automatically retrieves a specified number of triples in the one-hop subgraph where the entity resides; $\operatorname{Verbalize(\cdot)}$ is a function that transform the triple into its textual string, in this work, we use the linear verbalization (Baek et al., 2023); and $\operatorname{CONCAT(\cdot)}$ joins each verbalized triple statement with a separator.
+However, presenting information explicitly as structured triples may hinder effective knowledge retrieval by LLMs (Bian et al., 2024), since the structure of these subgraph triples substantially deviates from the textual data encountered during pre-training.
+Motivated by this limitation, we propose a second, more model-friendly approach, denoted as $D_{2}(e)$, which employs a language model to rephrase these verbalized triples into a narrative format that is more comprehensible to the LLM.
+Specifically, we employ an in-context learning approach, where a small set of carefully selected examples are prepended to the prompt to guide the LLM to generate high-quality, narrative-like descriptions.
+For brevity, Figure 3 illustrates only the core instruction of our prompt, omitting the in-context examples.
+This approach produces more refined and model-friendly entity descriptions, enabling the LLM to more effectively retrieve and leverage knowledge during the KGC task.
+By leveraging the subgraph entity description generator, we align the LLM’s internal understanding of entities with their counterparts in the KG, similar to an entity linking process.
+This approach addresses semantic ambiguity, achieving alignment comparable to fine-tuning while significantly reducing computational overhead.
+2.4. Representation Evaluation via Sliced Mutual Information
+To quantitatively assess the task-specific information encoded in the frozen-LLM representations, we adopt sliced mutual information (SMI), which scales more effectively to high-dimensional data compared to mutual information (Goldfeld and Greenewald, 2021; Wongso et al., 2023).
+Let the random variable $X$ denote the representations extracted from the intermediate layers, where $x_{k}=Rep_{j}^{l}(s_{k})\in\mathbb{R}^{d}$ corresponds to the representation of the $k$-th sample $s_{k}$.
+Let the random variable $Y$ denote the corresponding discrete KGC labels.
+Since labels $Y$ are discrete, the SMI is defined as the expectation of mutual information between the labels $Y$ and one-dimensional random projections of $X$, where these projections are sampled uniformly at random from the $d$-dimensional unit sphere (denoted $\mathbb{S}^{d-1}$).
+We estimate SMI via Monte-Carlo sampling combined with a scalar MI estimator (Kraskov et al., 2004):
+$\widehat{SMI}(X;Y)\;=\;\frac{1}{m}\sum_{i=1}^{m}\;\hat{I}\big{(}\{\theta_{i}^{\top}x_{k}\}_{k=1}^{n};\{y_{k}\}_{k=1}^{n}\big{)},\quad\theta_{i}\sim\text{Unif}(\mathbb{S}^{d-1}).$
+[…]
+3. Experiments
+In the experiments, our objective is to answer the following research questions (RQs):
+  • RQ1: Can LLMs with our method effectively achieve performance on par with their fine-tuned counterparts?
+  • RQ2: How does our method perform across different language models and classification models?
+  • RQ3: What are the data efficiency and computation efficiency of our proposed method for KGC?
+3.1. Datasets
+Dataset $|\mathbf{E}|$ $|\mathbf{R}|$ # Train # Valid # Test
+FB13 75,043 13 316,232 5,908 23,733
+WN11 38,696 11 112,581 2,609 10,544
+FB15K-237N 13,104 93 87,282 14,082 16,452
+WN18RR 40,943 11 86,835 6,068 6,268
+UMLS 135 46 5,216 1,304 1,322
+YAGO3-10 123,182 37 1,079,040 5,000 5,000
+Table 1. Statistical information of datasets.
+In our experiments, we employ five widely-used KG datasets for triple classification: FB13 (Socher et al., 2013), WN11 (Socher et al., 2013), FB15K-237N (Lv et al., 2022), WN18RR (Dettmers et al., 2018), and UMLS (Yao et al., 2019).
+For relation prediction, we adopt the YAGO3-10 dataset (Dettmers et al., 2018), and for entity prediction, we use WN18RR.
+Specifically, WN11 and WN18RR are subsets of WordNet, while FB13 and FB15K-237 (Toutanova and Chen, 2015) are subsets of Freebase.
+The FB15K-237N dataset is a modified version of FB15K-237.
+Additionally, UMLS is a medical semantic network comprising semantic types (entities) and semantic relations.
+To address incomplete or low-quality entity descriptions in some KGs, we propose two systematic approaches to generate entity descriptions from local subgraph structures, as detailed in Section 2.3 (hereafter referred to as generated entity descriptions).
+Specifically, for the first approach in Section 2.3, we construct entity descriptions by concatenating ten randomly sampled triples from each entity’s one-hop subgraph, employing linear verbalization (Baek et al., 2023); we denote these as generated descriptions (Tri).
+For the second approach, we further utilize GPT-3.5-turbo to rephrase these concatenated triples into coherent, model-friendly descriptions via in-context learning; we denote these as generated descriptions (GPT).
+To assess the quality and effectiveness of our generated entity descriptions, we also utilize existing entity descriptions (provided by the original datasets or external authoritative sources) as reference baselines, hereafter referred to as non-generated descriptions.
+Specifically, for FB13, entity descriptions are obtained from corresponding Wikipedia; for FB15K-237N, we adopt the descriptions provided by (Xie et al., 2016); for WN18RR, we utilize synset definitions from (Yao et al., 2019); and for UMLS, we employ the entity descriptions included in the original dataset.
 Method Pred Samples FB13 WN11 15K237N WN18RR UMLS
@@
 LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
-LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.934 0.862
-LLAMA-DES2 (Tri) H-MLP 3k/10k 0.842 0.755 0.671 0.779 0.782
-LLAMA-DES2 (GPT) H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
-LLAMA-DES-H H-MLP ALL 0.920 - 0.753 0.943 0.863
+LLAMA-MLP-DES (NG) H-MLP 3k/10k 0.901 - 0.738 0.934 0.862
+LLAMA-MLP-DES (Tri) H-MLP 3k/10k 0.842 0.755 0.671 0.779 0.782
+LLAMA-MLP-DES (GPT) H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
+LLAMA-MLP-DES (NG) H-MLP ALL 0.920 - 0.753 0.943 0.863
 Table 2.
@@
 ‘Samples’ indicates training sample counts, with $3k/10k$ indicating $10k$ samples for the UMLS dataset and $3k$ for others.
-LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimulus generation; LLAMA-DES-H combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-DES2 (Tri)/(GPT) uses first/second generated descriptions with $PT_{2}$ for stimulus generation.
+LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimulus generation; LLAMA-MLP-DES (NG) combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-MLP-DES (Tri)/(GPT) uses generated descriptions (Tri)/(GPT) with $PT_{2}$ for stimulus generation.
 Top three results per dataset are bolded.
-3. Experiments
-In the experiments, our objective is to answer the following research questions (RQs):
-  • RQ1: Can LLMs with our method effectively achieve performance on par with their fine-tuned counterparts?
-  • RQ2: How does our method perform across different language models and classification models?
-  • RQ3: What are the data efficiency and computation efficiency of our proposed method for KGC?
-3.1. Datasets
-Dataset $|\mathbf{E}|$ $|\mathbf{R}|$ Train # Valid # Test
-FB13 75,043 13 316,232 5,908 23,733
-WN11 38,696 11 112,581 2,609 10,544
-FB15K-237N 13,104 93 87,282 14,082 16,452
-WN18RR 40,943 11 86,835 6,068 6,268
-UMLS 135 46 5,216 1,304 1,322
-YAGO3-10 123,182 37 1,079,040 5,000 5,000
-Table 2. Statistical information of datasets.
-In our experiments, we employ five widely-used KG datasets for triple classification: FB13 (Socher et al., 2013), WN11 (Socher et al., 2013), FB15K-237N (Lv et al., 2022), WN18RR (Dettmers et al., 2018), and UMLS (Yao et al., 2019).
-For relation prediction, we employ the YAGO3-10 dataset (Dettmers et al., 2018), and for entity prediction, we use WN18RR for evaluation.
-For FB13, we utilize Wikipedia to construct entity descriptions.
-For FB15K-237N, the entity descriptions provided by (Xie et al., 2016) are used.
-Synset definitions from (Yao et al., 2019) served as entity descriptions for WN18RR.
-In the case of UMLS, entity descriptions provided by the dataset are used.
-In the following, we refer to these entity descriptions as non-generated entity descriptions.
 3.2. Experimental Settings
-To generate entity descriptions, for the first approach in Section 2.3, we concatenate ten randomly selected triples from the one-hop subgraph surrounding each entity to construct the descriptions.
-For the second approach, we employ GPT-3.5-turbo as the LLM.
-In the following text, we refer to these entity descriptions as generated entity descriptions (Tri)/(GPT).
-We compare our method with four structure information-based methods: TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), ComplEx (Trouillon et al., 2016) and RotatE (Sun et al., 2019), as well as seven additional information-based methods: KG-LLAMA (Yao et al., 2024), KG-BERT (Yao et al., 2019), KG-T5 (Saxena et al., 2022), LLaMA-7b, Alpaca-7B, LLaMA-7B-ICL and Alpaca-7B-ICL (Zhang et al., 2024).
-KG-LLAMA, KG-BERT and KGT5 fine-tune the language models using all the training samples.
-LLaMA-7B and Alpaca-7B rely on non-fine-tuned LLMs to directly perform predictions.
-LLaMA-7B-ICL and Alpaca-7B-ICL employ LLMs with in-context learning to make predictions (Zhang et al., 2024).
+Following the experimental settings in (Zhang et al., 2024), we compare our method with four widely-used structure information-based methods: TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), ComplEx (Trouillon et al., 2016) and RotatE (Sun et al., 2019).
+Additionally, we compare against seven additional information-based methods: KG-LLAMA (Yao et al., 2025), KG-BERT (Yao et al., 2019), KG-T5 (Saxena et al., 2022), LLaMA-7B, Alpaca-7B, LLaMA-7B-ICL and Alpaca-7B-ICL (Zhang et al., 2024).
+KG-LLAMA, KG-BERT and KG-T5 fine-tune the language models on the full training set.
+LLaMA-7B and Alpaca-7B use frozen LLMs directly for predictions.
+LLaMA-7B-ICL and Alpaca-7B-ICL perform predictions via in-context learning (Zhang et al., 2024).
+Since our method aims to fully exploit the capabilities of LLMs alone, we do not compare against hybrid methods such as KoPA (Zhang et al., 2024), which fine-tunes LLMs and additionally incorporates external structural methods (e.g., TransE).
+Such hybrid approaches incorporate additional structural information, and thus fall outside the scope of our current investigation.
 We give priority to using the results reported in the official baseline paper.
+For structure information-based methods, we reproduce the results with OpenKE (Han et al., 2018).
+For additional information-based methods, we reproduce results using official implementations.
 Please refer to Section 4.1 for detailed introductions of these baseline methods.
-For the negative sampling employed on the datasets, we utilize the same negative sampling technique as (Yao et al., 2019).
-For the intermediate layer selection, we identify the most suitable layer for the task by experimenting with the validation set.
-Each experiment is executed three times with random initializations, and the mean results are reported.
+Since existing KGs typically contain only positive triples, we follow standard practice (Yao et al., 2019) and generate negative samples by corrupting head or tail entities within positive triples for training.
+All training sets used in subsequent experiments thus contain equal number of positive and negative triples.
+For evaluation, we generate negative test triples equal in number to positive test triples on WN18RR, and UMLS datasets, while directly utilizing the original test splits containing balanced positive and negative triples for WN11, FB13 and FB15K-237N.
+For the intermediate layer selection, we determine the optimal layer via validation set performance.
+For the MLP classifier, we adopt a two-layer perceptron structure with PReLU activation and a softmax output layer.
+Specifically, we set the batch size to 64 and employ the AdamW optimizer with a learning rate of 3e-5.
+For the SVM classifier, we set the regularization parameter $C=0.5$ with a squared L2 penalty term and adopt the radial basis function (RBF) kernel.
+For the logistic regression classifier, we also set the regularization parameter $C=0.5$ and employ an L2 penalty term.
 3.3. Main Results (For RQ1)
-From Table 1, it can be observed that LLAMA-DES-H and LLAMA-DES2 (GPT), when using a partial training set, exhibit comparable performance to KG-LLAMA-7B, the latter being fine-tuned with all training triples.
-Moreover, when trained on subsets of equivalent size, LLAMA-DES-H and LLAMA-DES2 (GPT) markedly surpass KG-LLAMA-SAMPLED, the existing state-of-the-art approach that relies exclusively on LLaMA for KGC.
+As shown in Table 2, LLAMA-MLP-DES (NG) and LLAMA-MLP-DES (GPT), using only a limited subset (3k/10k) of the training samples, achieve comparable performance to KG-LLAMA-7B, which requires fine-tuning on the entire training set.
+Moreover, under equal-sized training subsets, LLAMA-MLP-DES (NG) and LLAMA-MLP-DES (GPT) significantly outperform KG-LLAMA-SAMPLED, the existing state-of-the-art approach that relies exclusively on LLaMA for KGC, with an average improvement of 17.15% and 18.16%.
 The experimental results presented in Table 2 also validate the effectiveness of the approaches proposed in Section 2.2 and Section 2.3.
-Specifically, LLAMA-DES-H and LLAMA-DES2 (GPT) outperform LLAMA-MLP, demonstrating the effectiveness of using subgraph entity description generator to improve the LLM’s understanding of entities in the KG, as proposed in Section 2.3.
+Specifically, LLAMA-MLP-DES (NG) and LLAMA-MLP-DES (GPT) outperform LLAMA-MLP, demonstrating the effectiveness of using subgraph entity description generator to improve the LLM’s understanding of entities in the KG, as proposed in Section 2.3.
 Furthermore, LLAMA-MLP demonstrates superior performance compared to LLAMA-7B and LLAMA-7B-ICL, affirming the effectiveness of stimulation method proposed in Section 2.2.
 Thus, the experimental results substantiate the effectiveness of our proposed methodology.
-Additionally, the non-fine-tuned LLM can also achieve similar KGC effects as that of fine-tuning the LLM.
-Notice that the performance on the UMLS dataset is special, where the use of generated descriptions hinders the model’s classification accuracy, and the use of non-generated entity descriptions is also worse than not using any entity descriptions.
-It is postulated that the underlying reason may lie in the highly specialized and domain-specific nature of entity names in the UMLS dataset, which are often precise medical terms with limited room for multiple interpretations.
+Additionally, our results clearly demonstrate that non-fine-tuned LLMs can achieve comparable KGC performance to their fine-tuned counterparts.
+Figure 4.
+SMI Values of Intermediate Layer Representations in the LLaMA Model Across Different Methods.
+W/O Description uses LLAMA-7B with $PT_{1}$ for stimulus generation; W/ Description uses LLAMA-7B with $PT_{2}$ for stimulus generation; SFT uses KG-LLAMA with $PT_{1}$ for supervised fine-tuning.
+Notably, the performance on the UMLS dataset exhibits a distinct pattern, where generated entity descriptions negatively impact classification accuracy, and non-generated descriptions yield worse results than not using descriptions as well.
+We hypothesize that this discrepancy arises from the highly specialized and domain-specific nature of entity names in the UMLS dataset, which are often precise medical terms with limited room for multiple interpretations.
 The entities in the knowledge graph have well-aligned with the internal entities of the LLM.
 Meanwhile, using entity descriptions may be affected by noisy description text, and descriptions generated by GPT-3.5-turbo may be influenced by hallucination, thereby affecting the classification effectiveness of the induced hidden states.
-Meanwhile, our method shows slightly lower performance on the WN11 and FB15K237N datasets compared to KG-LLAMA, which fine-tuned on the full training set.
+Meanwhile, our method shows slightly lower performance (3.8% and 2.2% lower, respectively) on the WN11 and FB15K237N datasets compared to KG-LLAMA, which was fine-tuned on the full training set.
 This discrepancy may stem from the fact that the pretrained LLM did not acquire knowledge of certain triples in the KG during its pretraining phase, resulting in incorrect predictions even when probing is conducted.
-In addition to the triple classification, we also validate the effectiveness of the proposed probing method on the relation prediction task as shown in Table 3.
-The experimental results similarly demonstrate that the probing method can achieve the nearly same performance as full training set fine-tuning, even when using only 0.6% of the training data.
-For entity prediction, we obtain the plausibility confidence scores of the triples through the classifier and rank them accordingly.
-Table 4 also displays similar results, showing that after using model-friendly descriptions, the LLAMA-DES2 (GPT) method improves the performance of LLAMA-MLP-7B and surpasses its fine-tuned counterparts.
-Figure 3.
-Differences in KGC performance across different layers of the LLAMA-DES2 (GPT) [left] and Mistral-DES2 (GPT) [right].
-The horizontal axis represents the position of the intermediate layers in LLaMA or Mistral, and the vertical axis represents the dataset.
 Method Pred Samples YAGO3-10
@@
 LLAMA-MLP-7B H-MLP 6996 0.5968
-LLAMA-DES2 (GPT) H-MLP 6996 0.6824
-LLAMA-DES2 (GPT) H-MLP ALL 0.7015
+LLAMA-MLP-DES (GPT) H-MLP 6996 0.6824
+LLAMA-MLP-DES (GPT) H-MLP ALL 0.7015
 Table 3. The experiment results (Hits@1) of relation prediction.
@@
 LLAMA-MLP-7B H-MLP 10k 0.2108
-LLAMA-DES2 (GPT) H-MLP 10k 0.2433
-LLAMA-DES2 (GPT) H-MLP ALL 0.2495
+LLAMA-MLP-DES (GPT) H-MLP 10k 0.2433
+LLAMA-MLP-DES (GPT) H-MLP ALL 0.2495
 Table 4. The experiment results (Hits@1) of entity prediction.
+In addition to the triple classification, we also validate the effectiveness of the proposed probing method on the relation prediction task as shown in Table 3.
+The experimental results similarly demonstrate that the probing method can achieve the nearly same performance as full training set fine-tuning, even when using only 0.6% of the training data.
+For entity prediction, we obtain the plausibility confidence scores of the triples through the classifier and rank them accordingly.
+Table 4 also displays similar results, showing that after using model-friendly descriptions, the LLAMA-MLP-DES (GPT) method improves the performance of LLAMA-MLP-7B and surpasses its fine-tuned counterparts.
+We further investigate the effectiveness of our method by evaluating the SMI values of representations extracted from intermediate layers, as illustrated in Figure 4.
+Notably, incorporating model-friendly entity descriptions (W/ Description) leads to substantial improvements in the SMI values compared to representations obtained without entity descriptions (W/O Description), yielding an average relative increase of 34.1% across the two datasets.
+This improvement substantiates that our subgraph entity description generator effectively aligns the LLM’s internal semantic comprehension with KGs’ entities, thereby yielding more precise intermediate representations that better predict triple plausibility.
+Moreover, representations enhanced by entity descriptions achieve SMI values comparable to those obtained from supervised fine-tuned (SFT) LLMs.
+This observation suggests that fine-tuning primarily aligns the LLM’s internal entity understanding with KG entities, rather than substantially augmenting the model’s intrinsic knowledge from the training set.
+These findings further underscore the feasibility and effectiveness of leveraging frozen LLMs for efficient and accurate KGC.
+These results further demonstrate the feasibility and effectiveness of leveraging frozen LLMs’ representations, combined with our proposed subgraph-based entity description generator, to achieve efficient and accurate KGC performance comparable to fine-tuned models.
+Figure 5.
+Differences in KGC performance across different layers of the LLAMA-MLP-DES (GPT) [left] and Mistral-MLP-DES (GPT) [right].
+The horizontal axis represents the position of the intermediate layers in LLaMA or Mistral, and the vertical axis represents the dataset.
 3.4. Ablation Study (For RQ2)
@@
 Intermediate layers closer to natural language text output generally exhibit better prediction performance compared to the lowest layers.
-This observation corroborates the notion posited by (Geva et al., 2020) that language models encode knowledge differentially across their Feed-Forward layers, with higher layers integrating information from lower layers.
-During pretraining, the lower layers may not have stored certain knowledge, and thus are unable to generate hidden states conducive to accurate triple classification.
+This observation aligns with the findings of (Geva et al., 2020), suggesting that language models encode knowledge differentially across Transformer layers, where higher layers integrate information from lower layers.
+During pretraining, the lower layers may not have stored certain knowledge, thus failing to produce hidden states informative for triple classification.
 Furthermore, the prediction performance of the layers closer to the natural language text output is generally worse than those in the middle layers.
-This phenomenon may be attributed to the hallucination issue inherent in LLaMA, wherein higher layers could potentially either omit or alter triple information, leading to a blurred distinction between positive and negative triple hidden states induced by stimuli (Jiang et al., 2024; Zou et al., 2023).
-Akin observations are discernible on the right side of Figure 3 for the Mistral model, which has a different backbone from LLaMA.
-This experimental results are consistent with the conclusions of (Jiang et al., 2024), which found that the MLP layers in the later stages have an inhibitory effect on the extracted knowledge, particularly contributing significantly to erroneous output decoding.
-As a result, the middle layers in their experiments also tend to outperform the later layers.
-Meanwhile, as observed in Table 6, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning with only a small sample size.
-Regarding classification models, a comparative analysis encompassing MLP, SVM, and Logistic Regression, as delineated in Table 5, is conducted.
-The result shows that MLP outperforms all three datasets compared to the other two classification models, while Logistic Regression also exhibits results close to MLP.
+This phenomenon may be attributed to the fact that intermediate layers appear to strike a balance between retaining sufficient information (avoiding overcompression) and discarding low-level noise, thus effectively preserving task-specific features while filtering out irrelevant details (Jiang et al., 2024; Zou et al., 2023).
+Similar trends can be observed on the right side of Figure 5 for the Mistral model, which employs an architecture distinct from LLaMA.
+Our empirical results are consistent with recent studies (Jiang et al., 2024; Orgad et al., 2024; Skean et al., [n.
+d.], 2025), highlighting the general advantage of intermediate layers over the final layers for downstream tasks.
+Meanwhile, as observed in Table 6, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning.
+We further compare different classifiers, including MLP, SVM, and Logistic Regression, with results summarized in Table 5.
+Results indicate that MLP consistently outperforms the other two classifiers across all three datasets, while Logistic Regression also exhibits results close to MLP.
 This underscores the efficacy of our stimulus methodology, whereby the model inherently encodes discriminative features for KGC tasks, making hidden states easily distinguishable in high-dimensional space and allowing straightforward classification boundaries to emerge without the need for complex, nonlinear decision functions.
@@
 LLAMA-MLP 0.855 0.872
-LLAMA-DES-H 0.901 -
-LLAMA-DES2 (GPT) 0.912 0.917
+LLAMA-MLP-DES (NG) 0.901 -
+LLAMA-MLP-DES (GPT) 0.912 0.917
 Mistral KG-MISTRAL-7B 0.891 0.962
 MISTRAL-MLP 0.846 0.901
-MISTRAL-DES-H 0.880 -
-MISTRAL-DES2 (GPT) 0.893 0.925
+MISTRAL-MLP-DES (NG) 0.880 -
+MISTRAL-MLP-DES (GPT) 0.893 0.925
 Gemma KG-GEMMA-7B 0.897 0.944
 GEMMA-MLP 0.806 0.893
-GEMMA-DES-H 0.891 -
-GEMMA-DES2 (GPT) 0.867 0.917
+GEMMA-MLP-DES (NG) 0.891 -
+GEMMA-MLP-DES (GPT) 0.867 0.917
+Qwen2.5 KG-Qwen-7B 0.913 0.961
+Qwen-MLP 0.864 0.908
+Qwen-MLP-DES (NG) 0.904 -
+Qwen-MLP-DES (GPT) 0.916 0.935
 Table 6.
-The variation in KGC prediction performance across different datasets using LLaMA, Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024) as base models (which have different backbones).
-Please refer to Table 1 for the number of training samples used by different methods.
+The variation in KGC prediction performance across different datasets using LLaMA, Mistral (Jiang et al., 2023), Gemma (Team et al., 2024) and Qwen2.5 (Qwen et al., 2025) as base models (which have different backbones).
 Method Procedure GPU Memory Time
@@
 3.5. Efficiency Study (For RQ3)
+Figure 6.
+The prediction performance of KG-LLAMA, KG-BERT, and LLAMA-MLP-DES (NG) on triple classification varies with the size of the training dataset.
+The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples.
 As illustrated in Figure 6, our method outperforms both KG-LLAMA and KG-BERT on the FB13 when trained with only 100 samples, as opposed to the 3000 samples used by the latter methods.
-For the UMLS, the performance of KG-LLAMA and KG-BERT with 3000 training samples is comparable to that of LLAMA-DES-H trained with only 500 samples.
+For the UMLS, the performance of KG-LLAMA and KG-BERT with 3000 training samples is comparable to that of LLAMA-MLP-DES (NG) trained with only 500 samples.
 Additionally, combining the results from Table 2, for the FB13 dataset, our method attains 97.2% of the performance achievable with the full training set, utilizing only 500 samples, which constitutes a mere 0.079% of the total training samples.
@@
 For the UMLS, with 9.58% of the training set, our method reaches 90.6% of the full training set performance.
-Our experimental results valid that the language model is capable of generating effective hidden states through our method, demonstrating its potential for data-efficient scenarios within the context of KGC tasks.
-In addition to data efficiency, we also evaluate the computational efficiency of our method.
-Table 7 shows that our method achieves significantly reduced resource usage compared to parameter-efficient fine-tuning.
-For the computation of GPU memory usage, we use ”torch.cuda.memory_allocated” to obtain the peak memory usage at each step.
-The KG-LLAMA method employs LoRA and 8-bit quantization technique for supervised fine-tuning.
-Compared to fine-tuned 7B LLMs, we boost the GPU memory efficiency to 188×188$\times$188 × during model training and speed up 26.11×26.$1$ overall.
+These results validate the capability of our method to effectively leverage the language model’s internal representations, highlighting its potential for data-efficient KGC scenarios.
+In addition to data efficiency, we investigate the computational efficiency of our method.
+Table 7 presents a detailed comparison between our approach with parameter-efficient fine-tuning, demonstrating substantial improvements in resource utilization.
+Specifically, GPU memory usage is measured using ‘torch.cuda.memory-allocated’ to record peak memory consumption per step.
+For a fair comparison, the KG-LLAMA baseline utilizes LoRA combined with 8-bit quantization techniques during supervised fine-tuning.
+As shown in Table 7, the time labeled as ‘Probing’ refers to the total inference duration required to extract representations for both training and test samples.
+Since this step involves only forward passes without gradient computations, we categorize it under the inference procedure.
+Compared to the fine-tuned 7B LLM baseline, our method improves GPU memory efficiency by $188\times$ during training and achieves an overall speedup of $26.11\times$.
 Notably, the training overhead of our method lies solely in the classifier, which is independent of the LLM size.
@@
 In these scenarios, the advantages of our method in terms of GPU memory usage and time consumption become even more pronounced.
-Figure 4.
-The prediction performance of KG-LLAMA, KG-BERT, and LLAMA-DES-H on triple classification varies with the size of the training dataset.
-The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples.
-Figure 5. Generalization across Datasets, LLAMA-DES-H. The horizontal axis represents the training set, and the vertical axis represents the test set.
-Figure 6. Performance of LLAMA-DES2 (GPT) versus the number of sampled triplets.
+Figure 7. Performance of LLAMA-MLP-DES (GPT) versus the number of sampled triplets.
+3.6. Additional Analysis
+We investigate how varying the number of sampled triples affects the performance of LLAMA-MLP-DES (GPT).
+As shown in Figure 7, increasing the number of sampled triples consistently improves accuracy.
+For example, on WN18RR, the accuracy increases from $86.6\%$ to $92.4\%$ as the number of triples grows from 1 to 20.
+This improvement highlights that richer subgraph information facilitates more precise entity descriptions, thus better aligning the LLM’s internal entity understanding with the KG semantics.
+Nonetheless, we also observe diminishing returns with increasing triples, indicating that additional subgraph information only marginally improves semantic alignment once sufficient context is captured.
+Meanwhile, as shown in Table 2, LLAMA-MLP-DES (Tri) consistently underperforms LLAMA-MLP-DES (GPT), exhibiting an average accuracy decrease of $10.5\%$ across all evaluated datasets.
+Surprisingly, LLAMA-MLP-DES (Tri) even falls short of LLAMA-MLP, which does not incorporate any form of entity description, resulting in an average accuracy decline of $6.24\%$.
+This suggests that naively concatenating verbalized triples with separators as entity descriptions does not facilitate—and may even impair—the LLM’s reasoning capabilities for KGC.
+A plausible explanation is that the structure of verbalized triples deviate significantly from the natural language distributions encountered by LLMs during pre-training, thereby introducing ambiguous stimulation information.
+In contrast, transforming these triples into a more natural, narrative-style format substantially alleviates this issue.
+Specifically, as demonstrated in Table 2, LLAMA-MLP-DES (GPT) achieves an average accuracy improvement of $3.96\%$ compared to LLAMA-MLP.
+This improvement underscores the effectiveness of generating model-friendly entity descriptions, enabling LLMs to retrieve and leverage relevant knowledge more precisely for KGC tasks.
+Additionally, to investigate the generalization capabilities of our approach, we conduct cross-dataset evaluation experiments where classifiers trained on one knowledge graph are tested on others.
+As shown in Figure 8, the classifier demonstrates strong generalization capabilities, particularly between datasets from similar knowledge sources.
+For instance, a classifier trained on FB13 achieves an accuracy of 0.75 when tested on FB15K237N, matching its in-domain performance.
+This substantial transfer performance can be attributed to the fact that both FB13 and FB15K237N are subsets of Freebase, sharing similar entity and relation structures.
+Classifiers trained on semantically distinct datasets, such as WN18RR (derived from WordNet) and UMLS, exhibit relatively weaker transfer performance when evaluated on Freebase-derived datasets.
+Nevertheless, these classifiers still achieve respectable performance, underscoring that representations probed from intermediate layers of LLMs encode generalizable and task-agnostic relational knowledge, enabling effective transfer across diverse knowledge graph contexts.
+Figure 8. Generalization across Datasets, LLAMA-MLP-DES (NG). The horizontal axis represents the training set, and the vertical axis represents the test set.
 4. Related Work
 4.1. Knowledge Graph Completion
-Knowledge graphs are typically characterized by their inherent incompleteness (Hogan et al., 2021).
-The prevailing methodologies in KGC can be categorized into two main classes: structure information-based methods and additional information-based methods (Shen et al., 2022b).
-Structure information-based methods leverage the inherent structural data present within knowledge graphs.
-For example, RESCAL (Nickel et al., 2011), DistMult (Yang et al., 2014), COMPGCN (Vashishth et al., 2019), and ConvE (Dettmers et al., 2018) primarily use scoring functions to evaluate the semantic similarities of entity or relation embeddings within a latent embedding space to infer missing information.
+Knowledge graphs are inherently incomplete, necessitating effective approaches for knowledge graph completion (Hogan et al., 2021).
+Prevailing KGC methodologies generally fall into two categories: structure information-based methods and additional information-based methods (Shen et al., 2022b).
+Structure information-based methods leverage the inherent structural patterns of knowledge graphs.
+Representative methods such as TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), COMPGCN (Vashishth et al., 2019), ConvE (Dettmers et al., 2018) and RotatE (Sun et al., 2019) primarily use scoring functions defined in a latent embedding space to infer missing information.
 Additional information-based methods incorporate a variety of supplementary data to enrich the knowledge graphs.
-The prominent examples are KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which represent entities, relations, and triples as textual sequences, turning KGC into a sequence classification problem.
-KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) employ encoder-decoder architecture language models to complete KGC in a generative manner.
-And (Li et al., 2024) enhances KGC with LLM by generating entity descriptions and designing loss functions for fine-tuning.
-More recently, KG-LLAMA (Yao et al., 2024) is the first to adopt a decoder-only LLM architecture for KGC by formulating the task as question answering through an instruction-tuning-based prompt, followed by parameter-efficient fine-tuning on LLaMA.
-Notably, KG-LLAMA leverages only LLaMA’s own capabilities without incorporating additional hybrid approaches, which aligns with the core principle of our proposed framework.
+Representative methods include KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which represent entities, relations, and triples as textual sequences, reformulating KGC as a sequence classification problem solvable via encoder-only language models.
+Meanwhile, KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) leverage encoder-decoder language models to perform KGC in a generative manner.
+Recently, decoder-only large language models have shown great promise in KGC tasks.
+KoPA (Zhang et al., 2024) integrates structural embeddings with LLMs to jointly leverage the structural information of KGs and their powerful contextual modeling capabilities.
+More recently, KG-LLAMA (Yao et al., 2025) frames KGC as an instruction-based question-answering task, applying parameter-efficient fine-tuning techniques to LLMs.
+Unlike hybrid approaches that combine LLMs with external structural embeddings, KG-LLAMA solely relies on the intrinsic capabilities of LLaMA, aligning closely with the core motivation of our proposed framework.
 4.2. Large Language Models
-Several studies have explored the mechanisms through which Large Language Models process, interpret, and generate factual content.
-As highlighted by (Saier et al., 2023), these models function as extensive repositories of knowledge, encapsulating a vast array of information about the world within their parameters.
-This insight into LLMs underscores their potential as powerful tools for a wide range of applications, including knowledge graph completion, by leveraging the rich informational content stored in their intricate network structures.
-(Ramrakhiyani et al., 2023) have conducted evaluations across multiple Large Language Models to assess their knowledge of world geography using probing techniques.
-Similarly, (Jin et al., 2024) used probing to measure the abilities of different LLM layers to learn various levels of concepts.
-In addition, (Liu et al., 2024) investigated whether a universal truthfulness hyperplane exists within LLMs by employing probing methods.
+Several studies have explored the mechanisms through which Large Language Models process, interpret, and generate factual knowledge.
+As demonstrated by (Saier et al., 2023), LLMs serve as implicit knowledge bases, encoding extensive world knowledge within their parameters.
+This intrinsic capability positions LLMs as promising tools for knowledge-intensive tasks such as knowledge graph completion, by effectively leveraging their embedded factual representations.
+To examine the extent and quality of factual knowledge encoded in LLMs, recent works have employed probing methods.
+For instance, (Ramrakhiyani et al., 2023) conducted evaluations across multiple LLMs to assess their knowledge of world geography using probing.
+Similarly, (Jin et al., 2024) analyzed the conceptual abstraction capabilities of different model layers, revealing hierarchical encoding patterns within the intermediate representations.
+Furthermore, (Liu et al., 2024) investigated the potential existence of a universal truthfulness hyperplane in LLMs via probing.
+Notably, recent probing analyses (Skean et al., 2025, [n.
+d.]) have indicated that intermediate layers often yield more informative representations for downstream tasks than the final layers.
+These findings align closely with our approach of leveraging intermediate representations from LLMs to achieve effective and efficient KGC.
 5. Conclusion